### v1

# Principal-Agent Multitasking: the Uniformity of Optimal Contracts and its Efficient Learning via Instrumental Regression

Abstract
This work studies the multitasking principal-agent problem. I first show a “uniformity” result. Specifically, when the tasks are perfect substitutes, and the agent’s cost function is homogeneous to a certain degree, then the optimal contract only depends on the marginal utility of each task and the degree of homogeneity. I then study a setting where the marginal utility of each task is unknown so that the optimal contract must be learned or estimated with observational data. I identify this problem as a regression problem with measurement errors and observe that this problem can be cast as an instrumental regression problem. The current works observe that both the contract and the repeated observations (when available) can act as valid instrumental variables, and propose using the generalized method of moments estimator to compute an approximately optimal contract from offline data. I also study an online setting and show how the optimal contract can be efficiently learned in an online fashion using the two estimators. Here the principal faces an exploration-exploitation tradeoff: she must experiment with new contracts and observe their outcome whilst at the same time ensuring her experimentations are not deviating too much from the optimal contract. This work shows when repeated observations are available and agents are sufficiently “diverse", the principal can achieve a very low $\widetilde{O}(d)$ cumulative utility loss, even with a “pure exploitation" algorithm.
1 Introduction
A firm’s owner (the principal) hires a manager (the agent) to manage the firm. The manager is responsible for several tasks, for example, managing personnel, maintaining a healthy work environment, planning and executing activities, etc. The manager can allocate a certain effort to each task, however, this effort is unobservable and unverifiable to the owner of the firm. Instead, the owner can observe a noisy signal, indicating the performance of each task. The principal can also observe the total revenue generated by the firm under the manager’s management, which will depend on the effort the manager puts into the tasks. The owner of the firm must incentivize the manager to exert effort through contracts. Since the exact effort is unobservable, the contract can only depend on the noisy signals associated with each task. Such is the basic model for the multitask principal-agent problem, which was first studied in the seminal work [1].
It often happens that the principal contracts with more than one agent. For example, in franchising scenarios, the principal is the franchiser and the agents are franchisees ([2]). The franchiser must design contracts to incentivize each franchisee to exert the appropriate amount of effort to generate revenue. Another example is e-commerce platforms, such as YouTube. YouTube is a video-sharing platform where content creators earn revenue through advertisements. According to this documentation ([3]), YouTube content creators earn 55% of the total advertisement revenue, while the remaining 45% goes to YouTube. While intuition might suggest that the principal should tailor her contracts when facing different agents to maximize her utility, in the above two scenarios, it is often observed that the contract is uniform across different agents. This leads to our first research question.
How can the uniformity of the contract terms be justified in (multitask) principal-agent problems?
As an answer to this question, I show that when the agent’s cost function is homogeneous to a certain degree and the principal’s utility takes a linear form across tasks (i.e., tasks are perfect substitutes), then the optimal contract is indeed uniform for different agents. In particular, the optimal contract only depends on the marginal utility of each task and the degree of homogeneity of the agent’s cost function. The implications of this result are practically very attractive. It implies that there is no need for the principal to discriminate against agents in an attempt to maximize her own utility, i.e., the optimal contract is fair. The result here follows closely in spirit to a result in the important work by [2]; the difference is that in their work they study a single task setting with double moral hazard while this work studies a multitask setting.
The above result shows that in the optimal contract, the principal should transfer more utility to the agent for more “important” tasks, which aligns with intuition. However, it may be the case that the exact marginal utility of each task is unknown. In this case, the principal must learn this value with observational data in order to design the optimal contract. This leads to our second research question.
How can the principal learn the optimal contract when the “importance” of each task is unknown?
Assuming again that the principal’s utility takes a linear form across tasks, estimating the marginal utilities of each task is essentially a linear regression problem, where the total revenue is the response variable, and the observed signals of each task are the covariates. However, since the observed signals are noisy measurements of the true effort, the regression problem has a measurement error problem ([4]). This work identifies the contract as a valid instrumental variable and proposes estimation and learning algorithms based on the generalized method of moments (GMM) estimator ([5]). Moreover, when repeated observations are available, the GMM estimator based on repeated observations achieves even faster convergence when the agents are “diverse”.
It seems important to point out that a problem setting very similar to the current work was studied in the (fairly) recent work [6]. Their work studied optimization algorithms in Stackelberg games from revealed preferences. As a special case, they gave a discussion on how their algorithm can be applied to the multitask principal-agent problem (though their work did not explicitly identify their problem as such). The problem of learning the optimal contract is also studied in this work, however, the approach is fundamentally different from theirs. This is because the current work also identified a “uniformity” result, which seemed to escape their analysis. The problem of estimating and learning the optimal contract in this work uses instrumental regression with the generalized method of moments method and is based on this “uniformity” result, whereas in their work the problem is solved via a two-stage optimization algorithm.
2 Related Work
Contract Theory
Contract theory studies a principal-agent problem where the principal must incentivize the agent to exert effort through contracts. Some early important works in the economics community include [7, 8, 1]. The work by [1] studies the multitask principal-agent problem, which serves as the starting point of the current work. Recently, computational aspects of contract design have been studied by many works in the computer science community. For example, [9, 10, 11] study combinatorial algorithms in contract design; [12, 13, 14] study learning algorithms for contract design.
Instrumental Regression and Measurement Error Models
Measurement error model is a well-studied problem in econometrics and statistics, for a textbook treatment the reader is referred to [4]. The instrumental regression and generalized method of moments (GMM) estimators (again for a textbook treatment see [5]) are popular methods to address endogeneity issues in models, including the measurement error model. Some recent works in computer science also notice endogeneity issues in learning/classification problems and apply the instrumental regression method in the context of strategic classification ([15]), online learning/bandits ([16]), reinforcement learning ([17]), etc.
Bandits and Exploration-Exploitation Tradeoff
Bandit problems are a type of online learning problem with partial feedback. Typically, bandit algorithms need to balance exploration (experimenting with unknown actions) and exploitation (choosing actions whose reward is estimated to be higher). Research on bandit problems is too broad to cover here, but for an overview see [18]. Recently, a number of papers have sought to understand the performance of the pure exploitation (i.e., greedy) algorithm in bandit problems. In particular, for linear bandits, the success of the greedy algorithm is explained under the framework of smoothed analysis [19, 20, 21].
3 Multitask Principal-Agent: Problem Description and the Uniformity Result
I study a multitasking principal-agent problem. The agent can exert effort in $d$ dimensions, where each dimension can be interpreted as a task. Denote effort of the agent by the vector $w\in(\mathbb{R}^{d})^{+}$, and assume that the agent incurs a private cost $c(w)\in\mathbb{R}^{+}$ when exerting effort $w$. The exact effort is unobservable and unverifiable to the principal, but a signal $x\in\mathbb{R}^{d}$ is generated and observed by the principal. The signal is assumed to satisfy
$\mathbb{E}[x|w]=w.$
The signal can be interpreted as a noisy measurement of the agent’s true effort.
The agent’s effort gives a private benefit $y$ to the principal, and it is assumed this benefit depends on the agents true efforts through the linear model:
$\mathbb{E}[y|w]=\langle\theta,w\rangle.$
Here $\theta\in(\mathbb{R}^{d})^{+}$. The value $\theta_{i}$ can be interpreted as the marginal utility of task $i$.
The principal incentivizes the agent to exert effort through a contract $\beta\in\mathbb{R}^{d}$. Under this contract, the monetary transfer to the agent is $\langle\beta,x\rangle$ when the realized signal is $x$. The agent then takes the action $w=w(\beta)$ which maximizes his expected utility under the contract $\beta$:
$w=\arg\max_{w^{\prime}}\langle\beta,w^{\prime}\rangle-c(w^{\prime}).$
Then, under the contract $\beta$, assuming the agent takes his best action $w$, the principal’s expected utility is
$\langle\theta,w\rangle-\langle\beta,w\rangle,$
i.e., the expected private benefit minus the expected payment to the agent. The optimal contract is the contract maximizing the principal’s expected utility.
Definition 1.
A function $c$ is homogeneous of degree $k$, if for any $\rho>0$ and $w\in(\mathbb{R}^{d})^{+}$, $c(\rho w)=\rho^{k}c(w)$.
The parameter $k$ is also known as the return to scale parameter.
Definition 2.
The cost function $c$ is said to satisfy the Inada conditions if the following holds.
  1. 1.
    $c$ is convex on its domain.
  2. 2.
    $\lim_{w_{i}\rightarrow 0}\partial c(w)/\partial w_{i}=0$.
  3. 3.
    $\lim_{w_{i}\rightarrow+\infty}\partial c(w)/\partial w_{i}=+\infty$.
The Inada condition is fairly standard. It essentially means the marginal cost goes from 00 to +$\infty$+ ∞ as $w_{i}$ goes from 00 to +$\infty$+ ∞. Throughout the following, the following assumption will be imposed on the cost function of the agent(s).
Assumption 1.
The cost function of the agent is homogeneous of degree $k$, and satisfies the Inada conditions.
3.1 The Uniformity Result
This subsection shows the optimal contract depends on the shape of $c(\cdot)$ only through the degree of homogeneity.
Theorem 1.
The principal’s optimal contract is $\beta=\theta/k$.
Proof.
When the posted contract is $\beta$, the agent’s expected utility when taking action $w$ is
$\langle\beta,w\rangle-c(w).$
The first-order condition gives:
$\beta=\nabla c(w).$
The first-order condition must bind under 1.
The principal’s expected utility, written as a function of $w$, is then
$\langle\theta,w\rangle-\langle\beta,w\rangle=\langle\theta,w\rangle-\langle\nabla c(w),w\rangle=\langle\theta,w\rangle-kc(w).$
Note the above is a concave function in $w$. The first-order condition gives
$\theta=k\nabla c(w).$
But also recall at the agent’s optimum $\beta=\nabla c(w)$. Hence $\beta=\theta/k$ is the contract that maximizes the principal’s utility. ∎
3.2 Discussion of the Result
The uniformity result is particularly attractive from a practical perspective. It implies that the principal can achieve her maximum utility and ensure fairness simultaneously. In particular, there is no need to discriminate against agents whose productivity may differ (as long as their return to scale parameter $k$ remains the same). This standardization is particularly relevant in industries like franchising or E-commerce, where the principal (the franchiser or the E-commerce platform) contracts with multiple agents (franchisees, E-commerce workers / content creators), and profit must be shared between the two parties. From a practical perspective, using uniform contracts streamlines the creation of contracts and also promotes trust between principals and agents.
As a special case, the results explain the widespread 50-50 split rule seen in many scenarios (e.g., [22]). Specifically, when the agents have a quadratic cost function ($k=2$), the results indicate that the 50-50 split is precisely the optimal linear contract which ensures maximum utility for the principal.
4 Estimating the Optimal Contract: an Instrumental Regression Approach
Figure 1: Causal relationship between variables. The agent’s response is shaded indicating it is unobserved by the principal. The dashed item is a second set of observed signals, corresponding the the scenario studied in Section 5 when repeated observations are available.
Algorithm 1 Use Contract as IV: Explore then Commit Algorithm
[…]
for $t=1,\dots,\tau$ do
     // If using basis vector for ‘exploration’:
     Set $\beta_{t}=\mathbf{e}_{i}$, here $i=\lceil t/\sqrt{T}\rceil$
     // If using small random perturbation for ‘exploration’:
     Set $\beta_{t}=(\theta_{0}+\rho_{t})/k$, here $\rho_{t}$ is a suitable random vector with variance $\sigma_{0}^{2}$ in each coordinate
     Observe the signal $x_{t}\in\mathbb{R}^{d}$ and the total revenue $y_{t}\in\mathbb{R}$
end for
Record the data so far into $B_{\tau},X_{\tau},Y_{\tau}$
Use the GMM estimator and compute the estimate :
$\hat{\theta}=(B_{\tau}^{\top}X_{\tau})^{-1}B_{\tau}Y_{\tau}.$
for $t=\tau+1,\dots,T$ do
     Set $\beta_{t}=\hat{\theta}/k$ as the contract
end for
I now study a setting where the parameter $\theta$ is unknown, i.e., the marginal utility of each task is unknown. Stated a bit informally, the principal is unsure about the exact “importance” of each task. For example, when the manager (the agent) manages a firm and participates in multiple tasks, the firm’s owner (the principal) may be unsure of the exact importance of each task and how much each task contributes to the firm’s total revenue. When the parameter $\theta$ is unknown, the principal must estimate or learn this parameter through observed data. By the previous result (Theorem 1), which states the optimal contract is $\theta/k$, finding an approximately optimal contract is essentially equivalent to obtaining an accurate estimate for $\theta$. To make matters a bit simpler, assume the return to scale parameter $k$ is known and treated as a constant.
The data consists of tuples of the form $(\beta_{t},w_{t},x_{t},y_{t})$, representing the contract, agent’s true effort, observed signal, and total production outcome respectively. In the following, the total production outcome $y_{t}$ will just be referred to as the total revenue for short. Note that the observed data by the principal only include $(\beta_{t},x_{t},y_{t})$, and the agent’s true effort $w_{t}$ is hidden.
Similar to the previous section, it is assumed the observed signals for each task is unbiased:
$\mathbb{E}[x_{t}|w_{t}]=w_{t}$
and that the total revenue satisfies a linear relation:
$\mathbb{E}[y_{t}|w_{t}]=\langle\theta,w_{t}\rangle.$
Fix any agent’s type (i.e., a cost function that satisfies 1). Define $u(\beta)$ as the expected utility when posting contract $\beta$. The principal is interested in finding a contract $\beta$ that minimizes the utility loss, which is the expected utility difference between posting the optimal contract $\theta/k$ and the contract $\beta$:
$u(\theta/k)-u(\beta).$
The following two assumptions are imposed, which allows us to relate the utility loss to the squared estimation error.
Assumption 2.
$c_{0}\leq\theta_{i}\leq C_{0}$ for $i\in[d]$.
The above assumption is natural and relatively mild, it states that there is some lower and upper-bound on the marginal utility of each task.
Assumption 3.
There exists some constant $c$ so that for any $\hat{\theta}\in[c_{0},C_{0}]^{n}$, the following holds
|$\absolutevalue{u(\theta/k)-u(\hat{\theta}/k)}\leq c\cdot\norm{\theta/k-\hat{\theta}/k}_{2}^{2}.$
The above assumption can essentially be seen as a smoothness assumption on the function $u$. Then, if the principal has an estimate $\hat{\theta}$ and the principal posts $\beta=\hat{\theta}/k$ as the contract, the utility loss can be bounded by the square estimation error (up to constant factors):
|$\absolutevalue{u(\theta/k)-u(\hat{\theta}/k)}\leq O\left(\norm{\theta-\hat{\theta}}_{2}^{2}\right).$
The above will hold for any agent type as long as the assumptions are met. Therefore, the following will mainly focus on obtaining upper bounds on the quantity $\norm{\theta-\hat{\theta}}_{2}^{2}$.
The problem of finding the optimal contract can either take place in an offline or online setting. In the offline setting, a sequence of data consisting of observations $(\beta_{t},x_{t},y_{t})$ is given to the principal, and the principal must obtain an accurate estimate of the true parameter $\theta$.
In the online setting, the principal interacts with agents in a sequential fashion consisting of $T$ rounds. In this scenario, for each round $t$, the principal must choose the contract $\beta_{t}$ based on historical data, then observe the signals $x_{t}$ and the total revenue $y_{t}$. In this online setting, the principal may be interested in minimizing the cumulative utility loss. A difference from the offline setting is that the principal needs to engage in some ‘experimentation’. Specifically, the principal may want to strategically choose some contracts $\beta$ in order to get a better estimate of the true parameter $\theta$; however, this may come at the cost of losing immediate revenue.
4.1 Offline Setting: Estimation with Instrumental Regression
The causal relationship between $(\beta_{t},w_{t},x_{t},y_{t})$ is summarized in Figure 1. The problem of estimating $\theta_{t}$ is essentially a (linear) regression problem. It can be seen that the dependence between the variables form a measurement error model, and it is well-known that directly regressing on $x_{t}$ as an attempt to learn $\theta$ will lead to a biased estimate. However, it can be observed that the contract $\beta_{t}$ acts as a valid instrumental variable. Specifically, it can be observed that
$\mathbb{E}[\beta_{t}(y_{t}-\langle\theta,x_{t}\rangle)]=\mathbf{0}\ (\in\mathbb{R}^{d}).$
Therefore an instrumental regression approach and the generalized method of moments estimator can be applied. When given a sequence of offline data, the generalized method of moments estimator takes the following form:
$\hat{\theta}_{T}=(B^{\top}_{T}X_{T})^{-1}B^{\top}_{T}Y_{T}.$
Here $B_{T}\in\mathbb{R}^{T\times d}$ is a matrix with rows representing the posted contracts each round; $X_{T}\in\mathbb{R}^{T\times d}$ is a matrix with rows representing the signals each round; $Y_{T}\in\mathbb{R}^{T}$ is a column vector representing the outcomes each round. Denote $\gamma_{t}=y_{t}-\langle\theta,x_{t}\rangle$, and collect values for $\gamma_{1:t-1}$ into the column vector $\Gamma_{t}$.
Proposition 1.
The estimation error
$\norm{\theta-\hat{\theta}_{T}}_{2}\leq\frac{\norm{B_{T}^{\top}\Gamma_{T}}}{\sigma_{\min}(B_{T}^{\top}X_{T})}.$
Proof.
$\hat{\theta}_{T}$ $=(B_{T}^{\top}X_{T})^{-1}B_{T}^{\top}Y_{T}$
$=(B_{T}^{\top}X_{T})^{-1}B_{T}^{\top}(X_{T}\theta+\Gamma_{T})$
$=\theta+(B_{T}^{\top}X_{T})^{-1}B_{T}^{\top}\Gamma_{T}.$
The proposition then follows. ∎
Remark 1.
By a standard concentration inequality, the numerator can be upper bounded as $O(\sqrt{d{T}})$ with high probability. The estimation error then depends crucially on the minimum singular value of $B_{T}^{\top}X_{T}$.
4.2 Online Setting: Learning the Optimal Contract
The estimator above gives a estimation error bound when offline observations are available. However, it does not directly give rise to a learning algorithm in the online setting, where the principal needs to choose the contract each round. To give some intuition, in this setting, the principal essentially faces an optimal design problem with an exploration-exploitation tradeoff. Specifically, the matrix $B_{t}$, consisting of contracts $\beta_{t}$, is the design, and the principal must choose this matrix such that the estimate $\hat{\theta}$ becomes accurate. At the same time, the principal must ensure that the chosen contracts $\beta_{t}$ are close to the true optimal contract $\theta/k$ in order to guarantee a small cumulative utility loss.
Lower bound on regret: $\Omega(d\sqrt{T})$
Consider a setting where each agent has cost function $c(w)=\frac{1}{2}\norm{w}_{2}^{2}$, and that this cost function is known to the principal. Then, the agent’s best response under the contract $\beta$ is exactly $w=\beta$. The principal’s expected utility when a contract $\beta$ is posted is then
$u(\beta)=\langle\theta,\beta\rangle-\norm{\beta}_{2}^{2}.$
The optimal contract is then $\beta=\theta/2$ (this is also implied by the previous uniformity result). The principal now essentially faces a zero-order optimization problem with a quadratic objective function. The contract $\beta$ is the decision variable, and the realized total revenue is the observed variable. Further, by a result in [23], the tight regret bound for this problem is $\Omega(d\sqrt{T})$.
Simple algorithms achieving $O(d\sqrt{T})$ cumulative utility loss
Perhaps a bit surprisingly, a very simple algorithm can achieve $O(d\sqrt{T})$ regret. In the first $d\sqrt{T}$ rounds, for each $i\in[d]$, the principal should set the contract as $\beta=\mathbf{e}_{i}$ for $\sqrt{T}$ rounds. The agents best response $w\in\mathbb{R}^{d}$ should only have a non-zero value in the $i$-th component $w_{i}$. Hence, the data in these rounds can be used to estimate $\theta_{i}$, and the estimation error after the initial $d\sqrt{T}$ rounds is:
|$\absolutevalue{\hat{\theta}_{i}-\theta_{i}}=O(1/{T^{1/4}}).$
In the remaining rounds, the principal can simply use the estimate after the initial $d\sqrt{T}$ rounds, and post the contract as $\beta=\hat{\theta}/k$. The regret in these rounds can be bounded as:
$T\cdot\norm{\hat{\theta}-\theta}_{2}^{2}=d\sqrt{T}.$
[…]
5 Improving the Estimation with Repeated Observations and Diversity
Algorithm 2 Use Repeated Observation as IV: Pure Exploitation with Diversity
Parameters include time horizon $T$, number of tasks $d$, minimum eigenvalue $\lambda_{0}$, failure probability $\delta$
Epoch $e$ contains |$\absolutevalue{\tau_{e}}=\max\{d\cdot 2^{e},8dC_{0}K_{0}\ln(d/\delta)/(c_{0}^{2}\lambda_{0}),4\sigma_{0}^{2}d^{2}\ln(d^{2}/\delta)/(c_{0}^{2}\lambda_{0})\}$ rounds
In the 0-th epoch with $d$ rounds, post $\beta=\mathbf{e}_{i}$ for each $i\in[d]$ once and record $\widetilde{X}_{1},X_{1},Y_{1}$
for epoch $e=1,2,\cdots,\lceil\log(T/d)\rceil$ do
     Compute $\hat{\theta}_{e}$ using the GMM estimator from data in previous epoch:
$\hat{\theta_{e}}=(\widetilde{X}_{e}^{\top}X_{e})^{-1}X_{e}Y_{e}$
     Post $\beta_{t}=\hat{\theta}_{t}/2$ for each round in this epoch
     Record the data from this epoch into $\widetilde{X}_{e+1},X_{e+1},Y_{e+1}$
end for
This section continues the discussion where the principal must learn $\theta$ using sequential observations. The previous section showed that using an instrumental regression approach where the contract is the instrumental variable, the principal must balance exploration (ensuring that $\sigma_{\min}(B_{t}^{\top}X_{t})$ grows as $t$ increases) and exploitation (ensuring $\beta_{t}$ is close to the true optimal contract). In this section, I show that, under some additional assumptions, the learning rate of the principal can be greatly improved. Informally, I show that when the agents are sufficiently ‘diverse’, and that when repeated observations are available, the principal can achieve a logarithmic regret, even if the principal is using a ‘pure exploitation algorithm’.
To state the results in a more concise manner, this section restricts attention to the case where the cost functions of the agents have diagonal quadratic forms. Specifically, the cost function of the agent is parameterized by a vector $\kappa\in(\mathbb{R}^{d})^{+}$, so that the cost function of the agent is $c(w)=\frac{1}{2}w^{\top}\mathtt{diag}(\kappa)^{-1}w$. Then, the agent’s best response to a contract $\beta$ is the maximizer of $\langle\beta,w\rangle-c(w)$, which is equal to
$w=\kappa\odot\beta.$
I assume that the agent cost function in each round, parameterized by $\kappa_{t}$, is drawn from an unknown distribution. The distribution satisfies the following.
Assumption 4.
The minimum eigenvalue of $\mathbb{E}_{\kappa\sim\mathcal{D}}[\kappa\kappa^{\top}]$ is lower bounded by $\lambda_{0}$. Further, for any $\kappa$ in the support of $\mathcal{D}$, $\kappa\leq K_{0}$.
The key part in the above assumption is the minimum eigenvalue $\lambda_{0}$, it suggests that the agents’ are sufficiently “diverse" in their types. The second assumption is that repeated observations are available. I assume that for each agent, the principal can observe two sets of signals, denoted by $x_{t}$ and $\widetilde{x}_{t}$. The signals are $\sigma_{0}^{2}$-subgaussian vectors and satisfy the following:
$\mathbb{E}[x_{t}|w_{t}]=\mathbb{E}[\widetilde{x}_{t}|w_{t}]=w_{t}.$
Assumption 5.
The two sets of signals are assumed to be (conditionally) independent. Specifically:
$x_{t}\perp\widetilde{x}_{t}|w_{t}.$
The causal relationship is summarized in Figure 1 with the dashed items included. It can be observed that the second set of observations can serve as an instrumental variable for the other, specifically:
$\mathbb{E}[\widetilde{x}_{t}(y_{t}-\langle x_{t},\theta\rangle)]=\mathbf{0}\ (\in\mathbb{R}^{d}).$
Therefore, we can use the following GMM estimator to obtain an estimate $\hat{\theta}$:
$\hat{\theta}_{t}=(\widetilde{X}_{t}^{\top}X_{t})^{-1}\widetilde{X}_{t}^{\top}Y_{t}.$
The algorithm will simply use the estimator above, without regard to ‘exploration’. In the implementation, the algorithm proceeds in epochs with epochs doubling in length to save computation. In each epoch $e$, the principal sets the contract using the estimation $\hat{\theta}_{e}$ according to the above formula using data from the previous epochs. See Algorithm 2 for details.
Different from the algorithms in the previous section, the algorithm does not face the exploration-exploitation tradeoff, as the ‘diversity’ of the agents already acts as a form of exploration.
Theorem 2.
With probability $1-1/T$, the cumulative regret of the algorithm is $\widetilde{O}(d/\lambda_{0}^{2})$.
Proof Sketch. .
Consider an epoch $e$ with length |$\absolutevalue{\tau_{e}}$ |. The minimum singular value of $(\widetilde{X}_{e}^{\top}X_{e})$ is on the order of $\Omega(\absolutevalue{\tau_{e}}\lambda_{0})$. At the same time, the term $\norm{\widetilde{X}_{e}^{\top}Y_{e}}$ can be upper bounded by $\widetilde{O}(\sqrt{d\absolutevalue{\tau_{e}}})$. Hence, for each round in the epoch, the squared estimation error can be upper bounded by:
$\norm{\hat{\theta}_{e}-\theta}_{2}^{2}\leq\widetilde{O}(d/(\absolutevalue{\tau_{e}}\lambda_{0}^{2})).$
Therefore, the regret in each epoch is on the order of $\widetilde{O}(d/\lambda_{0}^{2})$. There are a total of $O(\log T)$ epochs, therefore the total regret is $\widetilde{O}(d/\lambda_{0}^{2})$. ∎
Remark 2.
At a technical level, the analysis is a combination of the convergence analysis of GMM estimators (e.g. Chapter 18 in [5]) and smoothed analysis in bandit problems ([19, 20, 21]).
Remark 3.
Interestingly, the “fast" convergence also holds under a different notion of ‘diversity’. Instead of assuming agents’ cost functions are diverse, consider a setting where agents do not necessarily best respond, and their response $w_{t}$ may be a random vector with some variance lower bounded from below. In this setting, the minimum singular value $\widetilde{X}_{t}^{\top}X_{t}$ still grows linearly with $t$, and the “fast" convergence rate is achieved.
6 Numerical Experiments
Figure 2: (a): Total loss using the Contract as IV, and exploration is achieved through small random perturbations; (b): Total loss using the repeated observations as IV and when the “diversity" condition is satisfied.
This section describes numerical simulations that test the performance of the learning algorithms based on the GMM estimators. In the experiments, the number of tasks is chosen as $d=5$, the total number of rounds is chosen as $T=10^{6}$, and the true parameter $\theta$ is chosen so that each entry is between $[1,3]$. Experiments are ran on a personal laptop.
The experiments have two settings. In the first setting, the contract is used as the instrumental variable, and small random perturbations are added as a means to achieve exploration in the exploration phase (Algorithm 1). The small perturbations are chosen as truncated Gaussian vectors so that the posted contract has components within $[1,3]$. The agents have cost function $c(w)=\norm{w}_{2}^{2}$. The cumulative utility loss of a single run is plotted in Figure 2(a). It can be seen the loss grows quickly in the warm start phase (which consists of $d\sqrt{T}\approx 5\times 10^{3}$ rounds), and then the regret grows slowly in the remaining rounds. The experiments are run for 20 trials, and the average cumulative loss in the 20 trials is 4134.7, roughly on the order of $O(d\sqrt{T})$.
In the second setting, repeated observations are used as the instrumental variable (Algorithm 2). The agent’s cost function is chosen as a diagonal quadratic function with form
$c(w)=\sum_{i=1}^{d}\frac{w_{i}^{2}}{2k_{i}}.$
The parameter $k_{i}$ is sampled from the uniform distribution on $[1,2]$ so that the ‘diversity’ condition is satisfied. The cumulative loss of a single run is plotted in Figure 2(b). The loss grows quickly in the 1st epoch (consisting of roughly $\widetilde{O}(d)$ rounds) and then grows very slowly in the remaining rounds. The experiment is repeated for 20 trials, and the average cumulative loss is 89.6 (recall the theoretical limit is $\widetilde{O}(d)$).
7 Conclusion and Limitation
This work studied the multitask principal-agent problem. I first present a result justifying the use of uniform contracts when the principal contracts with heterogeneous agents. Then, I study the problem of learning the optimal contract and identify the problem as a measurement error problem, and propose using instrumental regression with the generalized method of moments estimator to estimate and learn the optimal contract. A limitation of the current work is that in the “diversity” condition in Section 5, the agents are limited to have quadratic cost functions. I believe this can be extended to more general cost functions (beyond quadratic functions and beyond $k=2$). Specifically, the “diversity” condition will hold whenever there are some variations in the agent’s “abilities" in each task. I will leave a more formal treatment to future work.
Appendix A Proof of Theorem 2
Recall the estimator in each epoch
$\hat{\theta}_{e}=\left(\widetilde{X}_{e}^{\top}X_{e}\right)^{-1}\widetilde{X}_{e}Y_{e}.$
Denote $\gamma_{t}=y_{t}-\langle x_{t},\theta\rangle$, and collect $\gamma_{t}$ into the column vector $\Gamma_{e}$.
Proposition 2.
$\norm{\hat{\theta}_{e}-\theta}_{2}\leq\frac{\norm{\widetilde{X}_{e}\Gamma_{e}}}{\sigma_{\min}(\widetilde{X}_{e}^{\top}X_{e})}$.
To bound the estimation error, the following will lower bound $\sigma_{\min}(\widetilde{X}_{e}^{\top}X_{e})$ and upper bound $\norm{\widetilde{X}_{e}\Gamma_{e}}$ separately.
[…]
Lemma 1 (Matrix Chernoff, Tropp).
[…]
$\lambda_{\min}X_{k}\geq 0,\lambda_{\max}X_{k}\leq R.$
Define
$\mu_{\min}:=\lambda_{\min}(\sum_{k}\mathbb{E}X_{k}),$
$\mu_{\max}:=\lambda_{\max}(\sum_{k}\mathbb{E}X_{k}).$
Then
$\Pr[\lambda_{\min}(\sum_{k}X_{k})\leq(1-\delta)\mu_{\min}]\leq d\cdot\left[\frac{e^{-\delta}}{(1-\delta)^{1-\delta}}\right]^{\mu_{\min}/R}.$
The following Corollary may be easier to use than the above Lemma.
Corollary 1.
Using the same setup as above, then
$\Pr[\lambda_{\min}(\sum_{k}X_{k})\leq(1-\delta)\mu_{\min}]\leq d\exp(-\delta^{2}\mu_{\min}/(2R)).$
Lemma 2.
With high probability $\delta$, $\lambda_{\min}(\sum_{t\in\tau_{e}}w_{t}w_{t}^{\top})\geq\absolutevalue{\tau_{e}}c_{0}^{2}\lambda_{0}/2$.
Proof.
[…]
$\mathbb{E}[W_{t}]$ $=\mathbb{E}[w_{t}w_{t}^{\top}]$
$=\mathbb{E}[\mathtt{diag}(\kappa_{t})\theta_{t}\theta_{t}^{\top}\mathtt{diag}(\kappa_{t})]$
$=\mathbb{E}[\Theta_{t}\kappa_{t}\kappa_{t}^{\top}\Theta_{t}]$
$=\Theta_{t}\mathbb{E}[\kappa_{t}\kappa_{t}^{\top}]\Theta_{t}$
Recall $\lambda_{\min}(\mathbb{E}[\kappa_{t}\kappa_{t}^{\top}])\geq\lambda_{0}$. The following holds
$\lambda_{\min}(\mathbb{E}W_{t})\geq c_{0}^{2}\lambda_{0}.$
By Tropp’s lemma,
$\Pr[\lambda_{\min}(\sum_{k}W_{k})\leq\absolutevalue{\tau_{e}}c_{0}^{2}\lambda_{0}/2]$ $\leq d\cdot\exp(-\absolutevalue{\tau_{e}}c_{0}^{2}\lambda_{0}/(8dC_{0}K_{0}))$
$\leq\delta$
when
|$\absolutevalue{\tau_{e}}\geq\frac{8dC_{0}K_{0}}{c_{0}^{2}\lambda_{0}}\cdot\ln(d/\delta)$
Lemma 3.
With probability $1-2\delta$, for each epoch $e$ with length as specified in Algorithm 2, $\sigma_{\min}(\widetilde{X}_{e}^{\top}X_{e})\geq\absolutevalue{\tau_{e}}c_{0}^{2}\lambda_{0}/4$.
Proof.
In the following denote $\varepsilon_{t}=x_{t}-w_{t},\widetilde{\varepsilon}_{t}=\widetilde{x}_{t}-w_{t}$.
$\widetilde{X}_{e}^{\top}X_{e}$ $=\sum_{t\in\tau_{e}}x_{t}\widetilde{x}_{t}^{\top}$
$=\sum_{t\in\tau_{e}}(w_{t}+\varepsilon_{t})(w_{t}+\widetilde{\varepsilon}_{t})^{\top}$
$=\sum_{t\in\tau_{e}}w_{t}^{\top}w_{t}+\varepsilon_{t}^{\top}w_{t}+w_{t}\varepsilon_{t}+\varepsilon_{t}^{\top}\widetilde{\varepsilon}_{t}$
The previous lemma showed that with probability $1-\delta$,
$\lambda_{\min}(\sum_{t\in\tau_{e}}w_{t}^{\top}w_{t})\geq\absolutevalue{\tau_{e}}c_{0}^{2}\lambda_{0}/2.$
The remaining term
$\sum_{t\in\tau_{e}}\varepsilon_{t}w_{t}^{\top}+w_{t}\widetilde{\varepsilon}_{t}^{\top}+\varepsilon\widetilde{\varepsilon}^{\top}$
is a random matrix with mean $\mathbf{0}$. By a standard concentration inequality, the absolute value of each entry in the matrix can be upper-bounded as |$\sqrt{\absolutevalue{\tau_{e}}\sigma_{0}^{2}\ln(d^{2}/\delta)}$ with probability $\delta$. Then, since the minimum singular value must be upper bounded by the Frobenius norm:
$\sigma_{\min}\left(\sum_{t\in\tau_{e}}\varepsilon_{t}w_{t}^{\top}+w_{t}\widetilde{\varepsilon}_{t}^{\top}+\varepsilon\widetilde{\varepsilon}^{\top}\right)\leq\sqrt{\absolutevalue{\tau_{e}}\sigma_{0}^{2}d^{2}\ln(d^{2}/\delta)}.$
The conclusion is that with probability $1-2\delta$,
[…]
|$\geq\absolutevalue{\tau_{e}}c_{0}^{2}\lambda_{0}/4.$
[…]
add sometinghere, condition on the length of $\tau_{e}$ This finishes the proof. ∎
Lemma 4.
[…]
Proof.
Recall $\gamma_{t}=y_{t}-\langle x_{t},\theta\rangle$, and $\Gamma_{e}$ is the column vector collecting the variables $\gamma_{t}$. Note that $\gamma_{t}\widetilde{x}_{t}$ is a zero-mean, $\sigma_{0}C_{0}K_{0}$-subgaussian random vector and independent conditioned on all previous steps. The result then follows from a standard concentration inequality. ∎
Theorem 3.
With probability $1-1/T$, the total regret is $\widetilde{O}(d/\lambda_{0}^{2})$.
Proof.
By Lemma 4, Lemma 3, and Proposition 2, with probability $1-3\delta$, the regret in each epoch can be upper bounded as |$\absolutevalue{\tau_{e}}\cdot\widetilde{O}(d/(\absolutevalue{\tau_{e}}\lambda_{0}^{2}))=\widetilde{O}(d/\lambda_{0}^{2})$. Choose some appropriate $\delta$, e.g., $\delta<O(1/T^{2})$. Since there are at most $\log T$ epochs, the total regret will be $\widetilde{O}(d/\lambda_{0}^{2})$. ∎

### v2

@@
-Principal-Agent Multitasking: the Uniformity of Optimal Contracts and its Efficient Learning via Instrumental Regression
+Linear Contracts in Multitasking: Robustness, Uniformity, and Learning
 Abstract
-This work studies the multitasking principal-agent problem.
-I first show a “uniformity” result.
-Specifically, when the tasks are perfect substitutes, and the agent’s cost function is homogeneous to a certain degree, then the optimal contract only depends on the marginal utility of each task and the degree of homogeneity.
-I then study a setting where the marginal utility of each task is unknown so that the optimal contract must be learned or estimated with observational data.
-I identify this problem as a regression problem with measurement errors and observe that this problem can be cast as an instrumental regression problem.
-The current works observe that both the contract and the repeated observations (when available) can act as valid instrumental variables, and propose using the generalized method of moments estimator to compute an approximately optimal contract from offline data.
-I also study an online setting and show how the optimal contract can be efficiently learned in an online fashion using the two estimators.
-Here the principal faces an exploration-exploitation tradeoff: she must experiment with new contracts and observe their outcome whilst at the same time ensuring her experimentations are not deviating too much from the optimal contract.
-This work shows when repeated observations are available and agents are sufficiently “diverse", the principal can achieve a very low $\widetilde{O}(d)$ cumulative utility loss, even with a “pure exploitation" algorithm.
+In this work, we study the multitasking principal-agent problem.
+The agent performs several task for the principal, and the principal posts a contract incentivizing the agent to exert effort.
+The principal can observe a signal for each task, and the contract is a mapping from the space of possible signals to a payment.
+We study the special class of linear contracts from three perspectives: robustness, uniformity, and learning.
+Firstly, we show a robustness result: in an ambiguous setting when only first moment information is known, there is a linear contract maximizing the principal’s payoff in a worst-case scenario.
+Secondly, we show a uniformity result: when the agent’s cost function is homogeneous to a certain degree and the the principal’s utility takes a linear form across tasks, then the optimal contract depends on the agent’s cost function only through its homogeneuity degree.
+Thirdly, we study the problem of learning an optimal linear contract through observational data.
+We identify this as an measurement error model, and propose instrumental regression methods to estimate the optimal contract parameters in an offline setting, or to learn the optimal contract in an online setting.
 1 Introduction
-A firm’s owner (the principal) hires a manager (the agent) to manage the firm.
-The manager is responsible for several tasks, for example, managing personnel, maintaining a healthy work environment, planning and executing activities, etc.
-The manager can allocate a certain effort to each task, however, this effort is unobservable and unverifiable to the owner of the firm.
-Instead, the owner can observe a noisy signal, indicating the performance of each task.
-The principal can also observe the total revenue generated by the firm under the manager’s management, which will depend on the effort the manager puts into the tasks.
-The owner of the firm must incentivize the manager to exert effort through contracts.
-Since the exact effort is unobservable, the contract can only depend on the noisy signals associated with each task.
-Such is the basic model for the multitask principal-agent problem, which was first studied in the seminal work [1].
-It often happens that the principal contracts with more than one agent.
-For example, in franchising scenarios, the principal is the franchiser and the agents are franchisees ([2]).
-The franchiser must design contracts to incentivize each franchisee to exert the appropriate amount of effort to generate revenue.
-Another example is e-commerce platforms, such as YouTube.
-YouTube is a video-sharing platform where content creators earn revenue through advertisements.
-According to this documentation ([3]), YouTube content creators earn 55% of the total advertisement revenue, while the remaining 45% goes to YouTube.
-While intuition might suggest that the principal should tailor her contracts when facing different agents to maximize her utility, in the above two scenarios, it is often observed that the contract is uniform across different agents.
-This leads to our first research question.
-How can the uniformity of the contract terms be justified in (multitask) principal-agent problems?
-As an answer to this question, I show that when the agent’s cost function is homogeneous to a certain degree and the principal’s utility takes a linear form across tasks (i.e., tasks are perfect substitutes), then the optimal contract is indeed uniform for different agents.
-In particular, the optimal contract only depends on the marginal utility of each task and the degree of homogeneity of the agent’s cost function.
-The implications of this result are practically very attractive.
-It implies that there is no need for the principal to discriminate against agents in an attempt to maximize her own utility, i.e., the optimal contract is fair.
-The result here follows closely in spirit to a result in the important work by [2]; the difference is that in their work they study a single task setting with double moral hazard while this work studies a multitask setting.
-The above result shows that in the optimal contract, the principal should transfer more utility to the agent for more “important” tasks, which aligns with intuition.
-However, it may be the case that the exact marginal utility of each task is unknown.
-In this case, the principal must learn this value with observational data in order to design the optimal contract.
-This leads to our second research question.
+A firm’s owner (the principal) hires a manager (the agent) to perform multiple tasks, such as managing personnel, maintaining a healthy work environment, and executing strategic initiatives.
+The agent can allocate effort across these tasks, but this effort is unobservable and unverifiable to the principal.
+Instead, the principal receives noisy signals that indicate task performance and can observe the total revenue generated by the firm, which depends on the agent’s effort allocation.
+To incentivize the agent, the principal offers a contract that links compensation to these noisy signals.
+Since direct effort observation is impossible, the contract structure plays a crucial role in aligning incentives.
+This setting defines the multitask principal-agent problem, first studied in the seminal work [Holmstrom and Milgrom, 1991].
+A contract in a multitasking problem is a mapping from the signal space to a payment.
+In theory, such contracts can be highly complex.
+Even in the well-studied single-task setting, optimal contracts can exhibit significant complexity.
+However, empirical observations suggest that most contracts used in practice tend to be relatively simple.
+A particularly important class of simple contracts is linear contracts.
+Many recent works have examined linear contracts in single task settings from various perspectives, including robustness and approximation properties (e.g.
+[Dütting et al., 2019], [Carroll, 2015]).
+However, the role of linear contracts in multitask settings remains underexplored.
+In this work, we study linear contracts in multitask principal-agent problems from three key perspectives: robustness, uniformity, and learnability.
+Robustness.
+Recent works have studied the robustness of linear contracts and established their optimality in ambiguous or non-Bayesian worst-case settings (e.g.
+[Carroll, 2015], [Dütting et al., 2019]).
+In this work, we extend these findings to the multitask principal-agent setting and provide a robustness guarantee under ambiguity.
+Our research question can be summarized as follows:
+Does the linear contract exhibit strong robustness guarantees in the multitask principal-agent setting?
+We show that when the principal knows only the first moment of each agent’s action but lacks information about its full distribution, a linear contract remains optimal under the worst-case distribution.
+This result highlights the structural advantages of linear contracts in environments with uncertainty.
+Our findings are similar in spirit to the prior results in [Dütting et al., 2019], which demonstrate the robustness of linear contracts in single-task settings.
+However, the transition to multitask environments introduces additional complexities, requiring new proof ideas; the crux of our proof contains a triangulation result from convex geometry followed by an application of Kakutani’s fixed-point theorem.
+Uniformity.
+In many real-world scenarios, the principal engages with multiple agents under the same contractual framework.
+A prominent example is franchising, where the principal (franchiser) contracts with multiple franchisees to ensure consistent effort and revenue generation [Bhattacharyya and Lafontaine, 1995].
+Another example is e-commerce platforms such as YouTube, where content creators earn a share of advertising revenue.
+According to official documentation [Google, 2024], YouTube content creators receive 55% of the total advertisement revenue, while the remaining 45% is retained by the platform.
+While one might expect the principal to tailor contracts for different agents to maximize utility, it is often observed that contracts remain uniform across agents.
+This observation leads to our second research question.
+How can the uniformity of contract terms be justified in multitask principal-agent problems?
+To answer this question, we show that when the agent’s cost function exhibits a certain degree of homogeneity and the principal’s utility is linear across tasks, the optimal contract is necessarily uniform across agents.
+Specifically, the optimal contract depends on the agent’s cost function only through it’s the homogeneity degree.
+This result has strong practical implications: it suggests that the principal does not need to differentiate contracts to maximize utility, reinforcing fairness in contract design.
+Our findings are similar in spirit to the work of [Bhattacharyya and Lafontaine, 1995], who examine a single-task setting with double moral hazard.
+In contrast, our study focuses on multitask environments.
+Learning
+In the optimal contract, imaginably the principal should offer a higher incentives to tasks that are more “important”, i.e., tasks which have a higher marginal utility.
+However, the exact “importance” of each task may not always be known.
+In such cases, the principal must infer these values from observational data to design the optimal contract.
+This naturally leads to our third research question:
 How can the principal learn the optimal contract when the “importance” of each task is unknown?
-Assuming again that the principal’s utility takes a linear form across tasks, estimating the marginal utilities of each task is essentially a linear regression problem, where the total revenue is the response variable, and the observed signals of each task are the covariates.
-However, since the observed signals are noisy measurements of the true effort, the regression problem has a measurement error problem ([4]).
-This work identifies the contract as a valid instrumental variable and proposes estimation and learning algorithms based on the generalized method of moments (GMM) estimator ([5]).
-Moreover, when repeated observations are available, the GMM estimator based on repeated observations achieves even faster convergence when the agents are “diverse”.
-It seems important to point out that a problem setting very similar to the current work was studied in the (fairly) recent work [6].
-Their work studied optimization algorithms in Stackelberg games from revealed preferences.
-As a special case, they gave a discussion on how their algorithm can be applied to the multitask principal-agent problem (though their work did not explicitly identify their problem as such).
-The problem of learning the optimal contract is also studied in this work, however, the approach is fundamentally different from theirs.
-This is because the current work also identified a “uniformity” result, which seemed to escape their analysis.
-The problem of estimating and learning the optimal contract in this work uses instrumental regression with the generalized method of moments method and is based on this “uniformity” result, whereas in their work the problem is solved via a two-stage optimization algorithm.
-2 Related Work
+Assuming that the principal’s utility remains linear across tasks, estimating the marginal utilities reduces to a linear regression problem where her realized private benefit serves as the response variable and the observed signals for each task act as covariates.
+However, the observed signals are noisy representations of the agent’s true effort, leading to a measurement error model [Fuller, 2009].
+In classical regression, measurement error in covariates introduces bias, making naive least squares estimators inconsistent and unreliable.
+This bias arises because the noise in observed signals weakens the correlation between true effort and output, leading to attenuation effects where coefficient estimates are systematically biased toward zero.
+To address this issue, we leverage instrumental variable techniques, identifying the contract itself as a valid instrument.
+Specifically, we propose estimation and learning algorithms based on the generalized method of moments (GMM) estimator [Greene, 2008].
+The GMM approach corrects for the endogeneity introduced by measurement error by exploiting moment conditions that remain valid despite the noise.
+Furthermore, when repeated observations are available, they can serve as additional instrumental variables, enhancing estimation accuracy.
+The effectiveness of this method is particularly pronounced when agents exhibit sufficient diversity in their responses, facilitating a faster convergence rate to the true parameter.
+Specifically, we show that the principal can achieve a $\widetilde{O}(d)$ regret in utility loss, even adopting a pure exploitation-type algorithm.
+2 Additional Related Work
 Contract Theory
 Contract theory studies a principal-agent problem where the principal must incentivize the agent to exert effort through contracts.
-Some early important works in the economics community include [7, 8, 1].
-The work by [1] studies the multitask principal-agent problem, which serves as the starting point of the current work.
+Some early important works in the economics community include [Holmström, 1979, Holmstrom, 1982, Holmstrom and Milgrom, 1991].
+The work by [Holmstrom and Milgrom, 1991] studies the multitask principal-agent problem, which serves as the starting point of the current work.
 Recently, computational aspects of contract design have been studied by many works in the computer science community.
-For example, [9, 10, 11] study combinatorial algorithms in contract design; [12, 13, 14] study learning algorithms for contract design.
+For example, [Dütting et al., 2019, Duetting et al., 2024, Dütting et al., 2022] study combinatorial algorithms in contract design; [Ho et al., 2014, Zuo, 2024, Guruganesh et al., 2024] study learning algorithms for contract design.
 Instrumental Regression and Measurement Error Models
-Measurement error model is a well-studied problem in econometrics and statistics, for a textbook treatment the reader is referred to [4].
-The instrumental regression and generalized method of moments (GMM) estimators (again for a textbook treatment see [5]) are popular methods to address endogeneity issues in models, including the measurement error model.
-Some recent works in computer science also notice endogeneity issues in learning/classification problems and apply the instrumental regression method in the context of strategic classification ([15]), online learning/bandits ([16]), reinforcement learning ([17]), etc.
+Measurement error model is a well-studied problem in econometrics and statistics, for a textbook treatment the reader is referred to [Fuller, 2009].
+The instrumental regression and generalized method of moments (GMM) estimators (again for a textbook treatment see [Greene, 2008]) are popular methods to address endogeneity issues in models, including the measurement error model.
+Some recent works in computer science also notice endogeneity issues in learning/classification problems and apply the instrumental regression method in the context of strategic classification ([Harris et al., 2022]), online learning/bandits ([Della Vecchia and Basu, 2023]), reinforcement learning ([Yu et al., 2022]), etc.
 Bandits and Exploration-Exploitation Tradeoff
@@
 Typically, bandit algorithms need to balance exploration (experimenting with unknown actions) and exploitation (choosing actions whose reward is estimated to be higher).
-Research on bandit problems is too broad to cover here, but for an overview see [18].
+Research on bandit problems is too broad to cover here, but for an overview see [Lattimore and Szepesvári, 2020].
 Recently, a number of papers have sought to understand the performance of the pure exploitation (i.e., greedy) algorithm in bandit problems.
-In particular, for linear bandits, the success of the greedy algorithm is explained under the framework of smoothed analysis [19, 20, 21].
-3 Multitask Principal-Agent: Problem Description and the Uniformity Result
-I study a multitasking principal-agent problem.
-The agent can exert effort in $d$ dimensions, where each dimension can be interpreted as a task.
-Denote effort of the agent by the vector $w\in(\mathbb{R}^{d})^{+}$, and assume that the agent incurs a private cost $c(w)\in\mathbb{R}^{+}$ when exerting effort $w$.
-The exact effort is unobservable and unverifiable to the principal, but a signal $x\in\mathbb{R}^{d}$ is generated and observed by the principal.
-The signal is assumed to satisfy
-$\mathbb{E}[x|w]=w.$
+In particular, for linear bandits, the success of the greedy algorithm is explained under the framework of smoothed analysis [Bastani et al., 2021, Kannan et al., 2018, Sivakumar et al., 2022].
+3 The Model
+We study the multitasking principal-agent problem, where an agent exerts effort across $d$ dimensions, each representing a distinct task.
+Let the agent’s effort be denoted by the vector $a\in\mathcal{A}$, where $\mathcal{A}$ represents the set of feasible actions available to the agent.
+Depending on the setting, $\mathcal{A}$ may be $(\mathbb{R}^{+})^{d}$ or a subset thereof.
+For clarity, we specify assumptions on $\mathcal{A}$ as needed in each section.
+The agent incurs a private cost $c(a)\in\mathbb{R}^{+}$ when exerting the effort vector $a$.
+We will always assume $c(a)$ is strictly increasing, strictly convex, and twice differentiable.
+While the principal cannot directly observe or verify the exact effort level $a$, she receives a signal $x\in\mathbb{R}^{d}$, which serves as an imperfect but unbiased indicator of the agent’s effort:
+$\mathbb{E}[x|a]=a.$
 The signal can be interpreted as a noisy measurement of the agent’s true effort.
-The agent’s effort gives a private benefit $y$ to the principal, and it is assumed this benefit depends on the agents true efforts through the linear model:
-$\mathbb{E}[y|w]=\langle\theta,w\rangle.$
-Here $\theta\in(\mathbb{R}^{d})^{+}$. The value $\theta_{i}$ can be interpreted as the marginal utility of task $i$.
-The principal incentivizes the agent to exert effort through a contract $\beta\in\mathbb{R}^{d}$.
-Under this contract, the monetary transfer to the agent is $\langle\beta,x\rangle$ when the realized signal is $x$.
-The agent then takes the action $w=w(\beta)$ which maximizes his expected utility under the contract $\beta$:
-$w=\arg\max_{w^{\prime}}\langle\beta,w^{\prime}\rangle-c(w^{\prime}).$
-Then, under the contract $\beta$, assuming the agent takes his best action $w$, the principal’s expected utility is
-$\langle\theta,w\rangle-\langle\beta,w\rangle,$
-i.e., the expected private benefit minus the expected payment to the agent. The optimal contract is the contract maximizing the principal’s expected utility.
+The principal incentivizes the agent to exert effort through a contract.
+We focus primarily on linear contracts, parameterized by $\beta\in(\mathbb{R}^{+})^{d}$.
+Under such a contract, the agent receives a monetary transfer of $\langle\beta,x\rangle$ upon the realization of signal $x$.
+The agent, in turn, selects an action $a=a(\beta)$ that maximizes their expected utility given the contract terms:
+$a=\arg\max_{a^{\prime}}\langle\beta,a^{\prime}\rangle-c(a^{\prime}).$
+The agent’s effort $a$ gives a noisy private benefit $y(a)$ to the principal, and the principal’s expected payoff, or expected utility, is the expected private benefit minus the expected payment to the agent.
+In the expression below, letting $a$ be the best response to the contract $\beta$, the principal’s expected utility is then
+$u(\beta)=\mathbb{E}[y(a)]-\langle\beta,a\rangle.$
+To sum up, the interaction protocol can be summarized as follows.
+  1. 1.
+    The principal posts a contract $\beta\in\mathbb{R}^{d}$
+  2. 2.
+    The agent best responds with a private effort (action) $a\in\mathbb{R}$, and suffers private cost $c(a)$
+  3. 3.
+    A noisy signal $x$ is generated and that $\mathbb{E}[x|a]=a$
+  4. 4.
+    The agent’s expected utility is $\langle\beta,a\rangle-c(a)$
+  5. 5.
+    The principal’s expected utility is $\mathbb{E}[y(a)]-\langle\beta,a\rangle$.
+Affine Contracts
+In certain parts of the proof we will also consider affine contracts, which are linear contracts with the addition of an offset term.
+Under the affine contract $\gamma\in\mathbb{R}^{d+1}$, the payment to the agent (abusing notation) is $\gamma(x)=\gamma_{0}+\sum_{i\in[d]}\gamma_{i}x_{i}$.
+A Preliminary Lemma
+Before going into our technical results, we first give a preliminary lemma.
+Lemma 1 (Lemma 9 in [Roth et al., 2016]).
+Suppose the agent’s action set is downward-closed, meaning if $a\in\mathcal{A}$, then $\rho a\in\mathcal{A}$ for some $\rho\in(0,1)$.
+Among all non-negative linear or affine contracts that induce some non-zero action $a$, the linear contract $\beta=\nabla c(a)$ induces $a$ at least expected payment.
+4 Robustness in Ambiguous settings
+Here, we explore the robustness properties of linear contracts in an ambiguous setting.
+In particular, we show that linear contracts are worst-case optimal in environments where only the mean outcomes of each action are known.
+[…]
+The principal seeks to guarantee a worst-case payoff in an ambiguous setting where the exact distribution of each action is unknown.
+To formally define the worst-case payoff, we first introduce the concept of compatible distributions.
 Definition 1.
-A function $c$ is homogeneous of degree $k$, if for any $\rho>0$ and $w\in(\mathbb{R}^{d})^{+}$, $c(\rho w)=\rho^{k}c(w)$.
-The parameter $k$ is also known as the return to scale parameter.
-Definition 2.
-The cost function $c$ is said to satisfy the Inada conditions if the following holds.
-  1. 1.
-    $c$ is convex on its domain.
-  2. 2.
-    $\lim_{w_{i}\rightarrow 0}\partial c(w)/\partial w_{i}=0$.
-  3. 3.
-    $\lim_{w_{i}\rightarrow+\infty}\partial c(w)/\partial w_{i}=+\infty$.
-The Inada condition is fairly standard.
-It essentially means the marginal cost goes from 00 to +$\infty$+ ∞ as $w_{i}$ goes from 00 to +$\infty$+ ∞.
+[…]
+$\mathbb{E}_{x\sim F_{a}}[x].$
+A contract by a principal is a mapping from all possible outcomes to a payment, which we denote by $w:\mathcal{X}\rightarrow\mathbb{R}^{+}$.
+Note that the agent is protect by limited liability.
+The worst-case payoff for the principal in an ambiguous setting is defined as
+$u(w)=\min_{\mathcal{F}\text{ is compatible}}u(w;\mathcal{F})$ (1)
+Here $u(w;\mathcal{F})$ is the principal’s payoff using contract $w$ assuming the agent’s true distributions were $\mathcal{F}$ and that the agent best responds to the contract $w$.
+We show that there is a linear contract $\beta\in\mathbb{R}^{d}$ maximizing the above expression. Recall that under the linear contract $\beta$, the agent’s payment under signal $x$ is
+$\beta(x)=\sum_{i=1}^{d}\beta_{i}x_{i}.$
+Theorem 1.
+There exists a linear contract maximizing the principal’s worst-case objective as given in eq. 1.
+Remark 1.
+[…]
+4.1 Proof Sketch
+Let’s first focus on affine contracts.
+We show that there exists an affine contract that maximizes the principal’s worst-case objective.
+A key property of affine contracts is that under any family of compatible distributions over outcomes, the agent’s best response remains unchanged—so the principal’s payoff is independent of the particular distribution in that family.
+Then in the last step, we improve the affine contract to a linear contract.
+Our strategy consists of the following steps.
+From a Non-affine Contract to an Affine Contract.
+Given any (possibly non-affine) contract $w(\cdot)$, we will construct an affine contract $\phi$ and identify a corresponding family of compatible distributions $\mathcal{F}$ such that the principal’s payoff under $\phi$ is at least as high as under $w$ for the compatible distribution family $\mathcal{F}$.
+Since the payoff from an affine contract does not depend on the particular choice of distribution family $\mathcal{F}$, it follows that the worst-case objective achieved by some affine contract is at least as good as that achieved by $w(\cdot)$.
+A Family of Affine Contract via Upper Supporting Hyperplanes.
+To obtain the affine contract, we upper-bound $w(\cdot)$ by finding an affine function $\phi$ that satisfies
+$\phi(x)\geq w(x)\quad\text{for all }x.$
+Consider the lifted set
+$S=\{(x,w(x)):x\in\mathcal{X}\}.$
+For any hyperplane $H$ that supports $S$ from above (an upper supporting hyperplane), we can associate an affine contract $\phi(H)$ that satisfies the inequality above.
+In this way, $\phi(H)$ serves as an upper bound to the payment function $w$.
+“Self-Owned” Hyperplanes.
+We now introduce the notion of a self-owned hyperplane: a hyperplane $H$ is called self-owned if the induced action $a(H)$ under the affine contract $\phi(H)$ lies in the convex hull of its own contact set (after projected back to $\mathcal{X}$), that is,
+$a(H)\in\operatorname{conv}(P(H)).$
+Here we define
+$P(H)=\{x\in\mathcal{X}:(x,w(x))\in H\}.$
+When $H$ is self-owned, one may choose a compatible distribution where the support of taking action $a(H)$ is only supported on $P(H)$.
+Then, since $H$ touches the $S$, we know that the agent’s payment is identical under both $w$ and $\phi(H)$.
+This means that if the affine contract $\phi(H)$ induced action $a$, then under this compatible distribution the non-affine contract $w(\cdot)$ also induces action $a$, and that $\phi(H)$ induces $a$ at a no larger payment.
+Existence of a Self-Owned Hyperplane.
+A classical result in convex geometry (e.g.
+[Lee and Santos, 2017]) implies that the union of the convex hulls of $P(H)$ over all upper supporting hyperplanes $H$ covers the entire outcome space $\mathcal{X}$.
+In particular, for every action $a\in\mathcal{X}$ there exists some hyperplane $H$ such that $a\in\operatorname{conv}(P(H))$.
+Then, we can construct a set-valued mapping:
+$g:\mathcal{H}\to 2^{\mathcal{H}},$
+defined as
+[…]
+By applying Kakutani’s fixed point theorem ([Kakutani, 1941]), there exists a hyperplane $H^{*}$ such that
+$H^{*}\in g(H^{*}).$
+Then we have precisely found a self-owned hyperplane.
+Refining to a Linear Contract.
+Finally, we show that if the affine contract $\phi$ induces an action $a^{*}$, then there exists a linear contract (i.e., one with zero offset) that also induces $a^{*}$ while providing no greater payment to the agent.
+This completes the proof by establishing that within the class of contracts, a linear contract exists that maximizes the principal’s worst-case objective.
+By these steps, we conclude that for any arbitrary contract $w$, there is a linear contract that performs at least as well in the principal’s worst-case evaluation. This completes the proof sketch.
+4.2 An Example
+We illustrate the key ideas, in particular, the existence of a self-owned hyperplane, with a concrete example.
+Consider a setting with $d=2$ and the following non-affine (non-linear) contract $w$ defined on $\mathcal{X}=\{0,1\}^{2}$:
+$w(0,0)=0,\quad w(1,0)=2,\quad w(0,1)=2,\quad w(1,1)=3.$
+We lift the space $\mathcal{X}$ into $\mathbb{R}^{3}$ by appending the contract value as a third coordinate. In other words, we consider the set
+$S=\{(0,0,0),\ (1,0,2),\ (0,1,2),\ (1,1,3)\}.$
+If $w$ were affine, then $S$ would lie on a single hyperplane in $\mathbb{R}^{3}$.
+Since $w$ is not affine, $S$ forms a polytope with several facets.
+In particular, there are two upper facets of $S$ that yield upper supporting hyperplanes, each corresponding to an affine contract that upper bounds $w$.
+One such upper facet defines an affine contract $\phi_{1}$ with
+$\phi_{1}(0,0)=0,\quad\phi_{1}(1,0)=2,\quad\phi_{1}(0,1)=2,\quad\phi_{1}(1,1)=4.$
+This hyperplane touches $S$ at the points
+$X_{1}=\{(0,0),\ (1,0),\ (0,1)\}.$
+Similarly, another upper facet defines an affine contract $\phi_{2}$ given by
+$\phi_{2}(0,0)=1,\quad\phi_{2}(1,0)=2,\quad\phi_{2}(0,1)=2,\quad\phi_{2}(1,1)=3,$
+supported on the points
+$X_{2}=\{(1,0),\ (0,1),\ (1,1)\}.$
+Note the triangulation is clearly illustrated: $\mathcal{X}=\mathtt{conv}(X_{1})\cup\mathtt{conv}(X_{2})$.
+Besides the two upper facets, any convex interpolation between these two hyperplanes also yields a valid supporting hyperplane for $S$.
+Such interpolated hyperplanes are supported on the two endpoints $(1,0)$ and $(0,1)$.
+We now claim that there must exist at least one hyperplane among these—one that we call self-owned—for which the induced action lies within its own set of support points (after projecting back to $\mathcal{X}$).
+In other words, if $H$ is the hyperplane corresponding to an affine contract $\phi(H)$, define
+$P(H)=\{x\in\mathcal{X}\mid(x,w(x))\in H\}.$
+Then $H$ is self-owned if the action $a$ induced by $\phi(H)$ satisfies
+$a\in\operatorname{conv}(P(H)).$
+Suppose, for the sake of contradiction, that no self-owned hyperplane exists.
+Then the affine contract $\phi_{1}$ would induce an action falling in $X_{2}$ while $\phi_{2}$ would induce an action in $X_{1}$.
+However, the mapping from the hyperplane to the induced action is continuous (small changes in the hyperplane lead to small changes in the induced action).
+Therefore, as one continuously interpolates between $\phi_{1}$ and $\phi_{2}$, the induced action must eventually lie exactly in the convex combination (or “interval”) determined by $(1,0)$ and $(0,1)$.
+Note that here Kakutani’s fixed point theorem reduces to an intermediate value theorem.
+Hence, this guarantees the existence of a self-owned hyperplane.
+5 Uniformity of the Optimal Linear Contract
+In this section, we give a characterization of the optimal linear contract.
+Specifically, we present a uniformity result, which states that if the agent’s cost function exhibits a certain degree of homogeneity, then the optimal linear contract depends on the cost function only through its homogeneity degree.
+From this point onward, we assume that the agent’s action space is $\mathcal{A}=(\mathbb{R}^{+})^{d}$.
+We first introduce several standard definitions and outline the conditions imposed on the problem instance.
 Throughout the following, the following assumption will be imposed on the cost function of the agent(s).
 Assumption 1.
-The cost function of the agent is homogeneous of degree $k$, and satisfies the Inada conditions.
+The cost function of the agent is homogeneous of degree $k$. In other words, for any $\rho>0$ and $a\in(\mathbb{R}^{d})^{+}$, $c(\rho a)=\rho^{k}c(a)$.
+Remark 2.
+The parameter $k$ is also known as the return to scale parameter.
+We also assume the private benefit to the principal depends only on the agent’s true effort vector $a$ through a linear model.
+Assumption 2.
+There exists $\theta^{*}\in(\mathbb{R}^{d})^{+}$, such that
+$\mathbb{E}[y(a)]=\langle\theta^{*},a\rangle.$
+The value $\theta^{*}_{i}$ can be interpreted as the marginal utility of task $i$ to the principal.
 5.1 The Uniformity Result
-This subsection shows the optimal contract depends on the shape of $c(\cdot)$ only through the degree of homogeneity.
+Here we show the optimal linear contract depends on the shape of $c(\cdot)$ only through the degree of homogeneity.
+The proof is extremely simple, yet to the best of our knowledge, it has not appeared in prior work.
 Theorem 2.
-The principal’s optimal contract is $\beta=\theta/k$.
+The principal’s optimal contract is $\beta^{*}=\theta^{*}/k$.
 Proof.
-When the posted contract is $\beta$, the agent’s expected utility when taking action $w$ is
-$\langle\beta,w\rangle-c(w).$
-The first-order condition gives:
-$\beta=\nabla c(w).$
-The first-order condition must bind under 1.
-The principal’s expected utility, written as a function of $w$, is then
-$\langle\theta,w\rangle-\langle\beta,w\rangle=\langle\theta,w\rangle-\langle\nabla c(w),w\rangle=\langle\theta,w\rangle-kc(w).$
-Note the above is a concave function in $w$. The first-order condition gives
-$\theta=k\nabla c(w).$
-But also recall at the agent’s optimum $\beta=\nabla c(w)$. Hence $\beta=\theta/k$ is the contract that maximizes the principal’s utility. ∎
+Instead of finding the best optimal linear contract, let us find the best action of the agent to induce.
+Suppose the best action to induce were $a$.
+By Lemma 1, the best linear contract to induce $a$ is the contract $\beta=\nabla c(a)$.
+In other words, $\beta=\nabla c(a)$ induces $a$, and no other contracts induces $a$ at a lower payment.
+Then, the principal’s expected utility for inducing action $a$ is
+$\langle\theta^{*},a\rangle-\langle\beta,a\rangle$ $=\langle\theta^{*},a\rangle-\langle\nabla c(a),a\rangle$
+$=\langle\theta^{*},a\rangle-kc(a).$
+The above is a concave function of $a$, and the first-order condition gives:
+$\theta^{*}=k\nabla c(a).$
+But also recall we have $\beta=\nabla c(a)$. Hence at the optimal we must have $\beta^{*}=\theta^{*}/k.$
 5.2 Discussion of the Result
@@
 From a practical perspective, using uniform contracts streamlines the creation of contracts and also promotes trust between principals and agents.
-As a special case, the results explain the widespread 50-50 split rule seen in many scenarios (e.g., [22]).
-Specifically, when the agents have a quadratic cost function ($k=2$), the results indicate that the 50-50 split is precisely the optimal linear contract which ensures maximum utility for the principal.
-4 Estimating the Optimal Contract: an Instrumental Regression Approach
+A problem setting very similar to the current work was studied in the recent work [Roth et al., 2016].
+Their work studied optimization algorithms in Stackelberg games from revealed preferences.
+As a special case, they gave a discussion on how their algorithm can be applied to the multitask principal-agent problem (though their work did not explicitly identify their problem as such).
+However, the uniformity result which we presented here seemed to have somehow escaped their analysis.
+6 Estimating and Learning the Optimal Contract: an Instrumental Regression Approach
 Figure 1: Causal relationship between variables.
@@
 The dashed item is a second set of observed signals, corresponding the the scenario studied in Section 7 when repeated observations are available.
+This section builds upon the results in the previous section where we presented the uniformity result, but consider a setting where the parameter $\theta^{*}$ is unknown so that some estimation or learning is involved to obtain the optimal linear contract.
+In other words, the principal lacks precise knowledge of the relative ”importance” of each task.
+For instance, when a firm’s owner (the principal) hires a manager (the agent) to oversee multiple tasks, the owner may be uncertain about the exact contribution of each task to the firm’s overall performance.
+In such cases, the principal must estimate or learn $\theta^{*}$ using observational data.
+By the uniformity result (Theorem 2), which establishes that the optimal contract is given by $\beta^{*}=\theta^{*}/k$, identifying an approximately optimal contract reduces to obtaining an accurate estimate of $\theta^{*}$.
+We assume that the homogeneity degree parameter $k$ is known and treated as a constant.
+In other words, the return to scale parameter is a constant for different agent types.
+While one could also study the scenario where $\theta^{*}$ is known but $k$ is unknown, this case is less technically interesting.
+Since $k$ is just a single-dimensional parameter, it can be inferred using a some one-dimensional discretization-based algorithm.
+Thus, we focus on the case where $\theta^{*}$ is unknown while $k$ is known and fixed, which seems to be the more technically interesting problem.
+Note that while this bears some resemblance to the linear bandit problem, where seemingly $x$ represents the context and the private benefit $y$ represents the noisy realized reward, our problem is in fact quite different.
+This is because, in fact, the signal is a noisy realization of the true “context”, which is the agent’s private effort.
+Recall that $\mathbb{E}[y(a)]=\langle\theta^{*},a\rangle$, so $a$ is the true “context” vector in the linear bandit problem.
+Essentially, our problem faces a measurement error problem.
+We assume the principal has access to data in $T$ time periods.
+In time period $t$, the observed data consists of tuples of the form $(\beta_{t},x_{t},y_{t})$, representing the contract, observed signal, and private benefit respectively; the agent’s private effort $a_{t}$ is unobserved by the principal.
+The relationship is depicted in Figure 1.
+Similar to the previous section, it is assumed the observed signals for each task is unbiased:
+$\mathbb{E}[x_{t}|a_{t}]=a_{t},$
+and that the total revenue satisfies a linear relation:
+$\mathbb{E}[y_{t}(a_{t})]=\langle\theta^{*},a_{t}\rangle.$
+We make a mild assumption that coordinates of $\theta^{*}$ are bounded, so that there is some bounded set $\mathcal{B}$ in which the optimal contract $\beta^{*}$ belongs to.
+This is relatively mild.
+Even though the principal may not know the exact value of $\theta^{*}$, she may have some reasonable bound on the marginal utility of each task, i.e., on $\theta^{*}_{i}$.
+For example, if the principal knows that each $\theta_{i}$ will be bounded in the interval $[c_{0},C_{0}]$, then we can take $\mathcal{B}=[c_{0}/k,C_{0}/k]^{d}$.
+Without loss of generality we can perform normalization and assume the following holds.
+Assumption 3.
+There exists some set $\mathcal{B}\subset[0,1]^{d}$, such that $\beta^{*}\in\mathcal{B}$.
+Moreover, we assume the signals $x_{t}$ and realized private benefits $y_{t}$ are conditionally subgaussian.
+Assumption 4.
+Given the agent’s action $a_{t}$, the random signal $x_{t}$ and private benefit $y_{t}$ are $\sigma$-subgaussian.
+The problem of determining the optimal contract can be approached in either an offline or an online setting.
+In the offline setting, the principal is given a sequence of data consisting of observations $(\beta_{t},x_{t},y_{t})$ and must estimate the true parameter $\theta^{*}$ as accurately as possible.
+In the online setting, the principal interacts with agents sequentially over $T$ rounds.
+In each round $t$, the principal selects a contract $\beta_{t}$ based on past observations and then observes the corresponding signals $x_{t}$ and total revenue $y_{t}$.
+Unlike in the offline case, the principal now face a exploration vs exploitation tradeoff: while better estimates of $\theta^{*}$ improve contract effectiveness in the long run, acquiring these estimates often requires experimentation, which may lead to short-term revenue loss.
+6.1 Offline Setting: GMM Estimator
+The causal relationship between $(\beta_{t},a_{t},x_{t},y_{t})$ is summarized in Figure 1.
+The problem of estimating $\theta_{t}$ is essentially a (linear) regression problem.
+It can be seen that the dependence between the variables form a measurement error model, and it is well-known that directly regressing on $x_{t}$ as an attempt to learn $\theta^{*}$ will lead to a biased estimate.
+However, it can be observed that the contract $\beta_{t}$ acts as a valid instrumental variable.
+Specifically, it can be observed that the following moment condition is satisfied
+$\mathbb{E}[\beta_{t}(y_{t}-\langle\theta^{*},x_{t}\rangle)]=\mathbf{0}\ (\in\mathbb{R}^{d}).$
+Therefore an instrumental regression approach and the generalized method of moments estimator can be applied.
+When given a sequence of offline data, the generalized method of moments estimator takes the following form:
+$\hat{\theta}_{T}=(B^{\top}_{T}X_{T})^{-1}B^{\top}_{T}Y_{T}.$
+Here $B_{T}\in\mathbb{R}^{T\times d}$ is a matrix with rows representing the posted contracts each round; $X_{T}\in\mathbb{R}^{T\times d}$ is a matrix with rows representing the signals each round; $Y_{T}\in\mathbb{R}^{T}$ is a column vector representing the outcomes each round.
+Proposition 1.
+With probability at least $1-\delta$, the estimation error
+$\norm{\theta^{*}-\hat{\theta}_{T}}_{2}\leq\frac{\sqrt{dT\log(dT/\delta)}}{\sigma_{\min}(B_{T}^{\top}X_{T})}.$
+Proof.
+Denote $\gamma_{t}=y_{t}-\langle\theta^{*},x_{t}\rangle$, and collect values for $\gamma_{1:t-1}$ into the column vector $\Gamma_{t}$.
+$\hat{\theta}_{T}$ $=(B_{T}^{\top}X_{T})^{-1}B_{T}^{\top}Y_{T}$
+$=(B_{T}^{\top}X_{T})^{-1}B_{T}^{\top}(X_{T}\theta^{*}+\Gamma_{T})$
+$=\theta^{*}+(B_{T}^{\top}X_{T})^{-1}B_{T}^{\top}\Gamma_{T}.$
+Hence
+$\norm{\theta^{*}-\hat{\theta}_{T}}_{2}\leq\frac{\norm{B_{T}^{\top}\Gamma_{T}}}{\sigma_{\min}(B_{T}^{\top}X_{T})}.$
+By a standard concentration inequality (see Section A.1) the numerator can be upper bounded as $\sqrt{dT\log(dT/\delta)}$ with probability at least $1-\delta$, the proposition then follows. ∎
+6.2 Online Setting: Explore then Commit Learning Algorithm
 Algorithm 1 Use Contract as IV: Explore then Commit Algorithm
-[…]
+Number of exploration rounds $\tau=d\sqrt{T}$, an initial estimate $\theta_{0}$ for $\theta$
 for $t=1,\dots,\tau$ do
-     // If using basis vector for ‘exploration’:
      Set $\beta_{t}=\mathbf{e}_{i}$, here $i=\lceil t/\sqrt{T}\rceil$
-     // If using small random perturbation for ‘exploration’:
-     Set $\beta_{t}=(\theta_{0}+\rho_{t})/k$, here $\rho_{t}$ is a suitable random vector with variance $\sigma_{0}^{2}$ in each coordinate
      Observe the signal $x_{t}\in\mathbb{R}^{d}$ and the total revenue $y_{t}\in\mathbb{R}$
@@
 end for
-I now study a setting where the parameter $\theta$ is unknown, i.e., the marginal utility of each task is unknown.
-Stated a bit informally, the principal is unsure about the exact “importance” of each task.
-For example, when the manager (the agent) manages a firm and participates in multiple tasks, the firm’s owner (the principal) may be unsure of the exact importance of each task and how much each task contributes to the firm’s total revenue.
-When the parameter $\theta$ is unknown, the principal must estimate or learn this parameter through observed data.
-By the previous result (Theorem 1), which states the optimal contract is $\theta/k$, finding an approximately optimal contract is essentially equivalent to obtaining an accurate estimate for $\theta$.
-To make matters a bit simpler, assume the return to scale parameter $k$ is known and treated as a constant.
-The data consists of tuples of the form $(\beta_{t},w_{t},x_{t},y_{t})$, representing the contract, agent’s true effort, observed signal, and total production outcome respectively.
-In the following, the total production outcome $y_{t}$ will just be referred to as the total revenue for short.
-Note that the observed data by the principal only include $(\beta_{t},x_{t},y_{t})$, and the agent’s true effort $w_{t}$ is hidden.
-Similar to the previous section, it is assumed the observed signals for each task is unbiased:
-$\mathbb{E}[x_{t}|w_{t}]=w_{t}$
-and that the total revenue satisfies a linear relation:
-$\mathbb{E}[y_{t}|w_{t}]=\langle\theta,w_{t}\rangle.$
-Fix any agent’s type (i.e., a cost function that satisfies 1).
-Define $u(\beta)$ as the expected utility when posting contract $\beta$.
-The principal is interested in finding a contract $\beta$ that minimizes the utility loss, which is the expected utility difference between posting the optimal contract $\theta/k$ and the contract $\beta$:
-$u(\theta/k)-u(\beta).$
-The following two assumptions are imposed, which allows us to relate the utility loss to the squared estimation error.
-Assumption 2.
-$c_{0}\leq\theta_{i}\leq C_{0}$ for $i\in[d]$.
-The above assumption is natural and relatively mild, it states that there is some lower and upper-bound on the marginal utility of each task.
-Assumption 3.
-There exists some constant $c$ so that for any $\hat{\theta}\in[c_{0},C_{0}]^{n}$, the following holds
-|$\absolutevalue{u(\theta/k)-u(\hat{\theta}/k)}\leq c\cdot\norm{\theta/k-\hat{\theta}/k}_{2}^{2}.$
-The above assumption can essentially be seen as a smoothness assumption on the function $u$.
-Then, if the principal has an estimate $\hat{\theta}$ and the principal posts $\beta=\hat{\theta}/k$ as the contract, the utility loss can be bounded by the square estimation error (up to constant factors):
-|$\absolutevalue{u(\theta/k)-u(\hat{\theta}/k)}\leq O\left(\norm{\theta-\hat{\theta}}_{2}^{2}\right).$
-The above will hold for any agent type as long as the assumptions are met. Therefore, the following will mainly focus on obtaining upper bounds on the quantity $\norm{\theta-\hat{\theta}}_{2}^{2}$.
-The problem of finding the optimal contract can either take place in an offline or online setting.
-In the offline setting, a sequence of data consisting of observations $(\beta_{t},x_{t},y_{t})$ is given to the principal, and the principal must obtain an accurate estimate of the true parameter $\theta$.
-In the online setting, the principal interacts with agents in a sequential fashion consisting of $T$ rounds.
-In this scenario, for each round $t$, the principal must choose the contract $\beta_{t}$ based on historical data, then observe the signals $x_{t}$ and the total revenue $y_{t}$.
-In this online setting, the principal may be interested in minimizing the cumulative utility loss.
-A difference from the offline setting is that the principal needs to engage in some ‘experimentation’.
-Specifically, the principal may want to strategically choose some contracts $\beta$ in order to get a better estimate of the true parameter $\theta$; however, this may come at the cost of losing immediate revenue.
-4.1 Offline Setting: Estimation with Instrumental Regression
-The causal relationship between $(\beta_{t},w_{t},x_{t},y_{t})$ is summarized in Figure 1.
-The problem of estimating $\theta_{t}$ is essentially a (linear) regression problem.
-It can be seen that the dependence between the variables form a measurement error model, and it is well-known that directly regressing on $x_{t}$ as an attempt to learn $\theta$ will lead to a biased estimate.
-However, it can be observed that the contract $\beta_{t}$ acts as a valid instrumental variable.
-Specifically, it can be observed that
-$\mathbb{E}[\beta_{t}(y_{t}-\langle\theta,x_{t}\rangle)]=\mathbf{0}\ (\in\mathbb{R}^{d}).$
-Therefore an instrumental regression approach and the generalized method of moments estimator can be applied.
-When given a sequence of offline data, the generalized method of moments estimator takes the following form:
-$\hat{\theta}_{T}=(B^{\top}_{T}X_{T})^{-1}B^{\top}_{T}Y_{T}.$
-Here $B_{T}\in\mathbb{R}^{T\times d}$ is a matrix with rows representing the posted contracts each round; $X_{T}\in\mathbb{R}^{T\times d}$ is a matrix with rows representing the signals each round; $Y_{T}\in\mathbb{R}^{T}$ is a column vector representing the outcomes each round.
-Denote $\gamma_{t}=y_{t}-\langle\theta,x_{t}\rangle$, and collect values for $\gamma_{1:t-1}$ into the column vector $\Gamma_{t}$.
-Proposition 1.
-The estimation error
-$\norm{\theta-\hat{\theta}_{T}}_{2}\leq\frac{\norm{B_{T}^{\top}\Gamma_{T}}}{\sigma_{\min}(B_{T}^{\top}X_{T})}.$
-Proof.
-$\hat{\theta}_{T}$ $=(B_{T}^{\top}X_{T})^{-1}B_{T}^{\top}Y_{T}$
-$=(B_{T}^{\top}X_{T})^{-1}B_{T}^{\top}(X_{T}\theta+\Gamma_{T})$
-$=\theta+(B_{T}^{\top}X_{T})^{-1}B_{T}^{\top}\Gamma_{T}.$
-The proposition then follows. ∎
-Remark 1.
-By a standard concentration inequality, the numerator can be upper bounded as $O(\sqrt{d{T}})$ with high probability.
-The estimation error then depends crucially on the minimum singular value of $B_{T}^{\top}X_{T}$.
-4.2 Online Setting: Learning the Optimal Contract
-The estimator above gives a estimation error bound when offline observations are available.
+In the previous part, we developed an estimator based on GMM for estimating the parameter $\theta^{*}$, and by extension (in particular by the uniformity result), the optimal contract $\beta^{*}$.
+This applies to the setting when offline observations are available.
 However, it does not directly give rise to a learning algorithm in the online setting, where the principal needs to choose the contract each round.
@@
 Specifically, the matrix $B_{t}$, consisting of contracts $\beta_{t}$, is the design, and the principal must choose this matrix such that the estimate $\hat{\theta}$ becomes accurate.
-At the same time, the principal must ensure that the chosen contracts $\beta_{t}$ are close to the true optimal contract $\theta/k$ in order to guarantee a small cumulative utility loss.
-Lower bound on regret: $\Omega(d\sqrt{T})$
-Consider a setting where each agent has cost function $c(w)=\frac{1}{2}\norm{w}_{2}^{2}$, and that this cost function is known to the principal.
-Then, the agent’s best response under the contract $\beta$ is exactly $w=\beta$.
+At the same time, the principal must ensure that the chosen contracts $\beta_{t}$ are close to the true optimal contract $\beta^{*}=\theta^{*}/k$ in order to guarantee a small cumulative utility loss.
+We will be interested in bounding the following quantity, which is the square error of $\beta_{t}$ each round:
+$\mathtt{Reg}=\sum_{t=1}^{T}\norm{\beta^{*}-\beta_{t}}_{2}^{2}.$
+The above can serve as a proxy for the actual cumulative utility loss: $\sum_{t=1}^{T}\absolutevalue{u_{t}(\beta^{*})-u_{t}(\beta_{t})}$ |.
+In particular, as long as the function $u_{t}$ has a bounded second derivative at the global maximum $\beta^{*}$, the cumulative utility loss can be upper bounded by $O(\tt Reg)$.
+6.2.1 Upper Bounds
+Here we show that very simple explore-then-commit type algorithms can achieve $O(d\sqrt{T})$ regret.
+In the first $d\sqrt{T}$ rounds, for each $i\in[d]$, the principal should set the contract as $\beta=\mathbf{e}_{i}$ for $\sqrt{T}$ rounds.
+The agents best response $a\in\mathbb{R}^{d}$ should only have a non-zero value in the $i$-th component $a_{i}$.
+Hence, the data in these rounds can be used to estimate $\theta_{i}$.
+In the remaining rounds, the principal can simply use the estimate $\hat{\theta}$ after the initial $d\sqrt{T}$ rounds, and post the contract as $\beta=\hat{\theta}/k$.
+Proposition 2.
+There exists algorithm achieving regret $\widetilde{O}(d\sqrt{T})$ with probability $1-1/T$.
+The proof of the above can be found in Section A.2.
+Remark 3.
+[…]
+6.2.2 Lower Bound: $\Omega(d\sqrt{T})$
+Consider a setting where each agent has cost function $c(a)=\frac{1}{2}\norm{a}_{2}^{2}$, and that this cost function is known to the principal.
+Then, the agent’s best response under the contract $\beta$ is exactly $a=\beta$.
 The principal’s expected utility when a contract $\beta$ is posted is then
-$u(\beta)=\langle\theta,\beta\rangle-\norm{\beta}_{2}^{2}.$
-The optimal contract is then $\beta=\theta/2$ (this is also implied by the previous uniformity result).
+$u(\beta)=\langle\theta^{*},\beta\rangle-\norm{\beta}_{2}^{2}.$
+The optimal contract is then $\beta^{*}=\theta^{*}/2$ (this is also implied by the previous uniformity result, since the agent’s cost function is quadratic and homogeneous of degree $2$).
 The principal now essentially faces a zero-order optimization problem with a quadratic objective function.
 The contract $\beta$ is the decision variable, and the realized total revenue is the observed variable.
-Further, by a result in [23], the tight regret bound for this problem is $\Omega(d\sqrt{T})$.
-Simple algorithms achieving $O(d\sqrt{T})$ cumulative utility loss
-Perhaps a bit surprisingly, a very simple algorithm can achieve $O(d\sqrt{T})$ regret.
-In the first $d\sqrt{T}$ rounds, for each $i\in[d]$, the principal should set the contract as $\beta=\mathbf{e}_{i}$ for $\sqrt{T}$ rounds.
-The agents best response $w\in\mathbb{R}^{d}$ should only have a non-zero value in the $i$-th component $w_{i}$.
-Hence, the data in these rounds can be used to estimate $\theta_{i}$, and the estimation error after the initial $d\sqrt{T}$ rounds is:
-|$\absolutevalue{\hat{\theta}_{i}-\theta_{i}}=O(1/{T^{1/4}}).$
-In the remaining rounds, the principal can simply use the estimate after the initial $d\sqrt{T}$ rounds, and post the contract as $\beta=\hat{\theta}/k$. The regret in these rounds can be bounded as:
-$T\cdot\norm{\hat{\theta}-\theta}_{2}^{2}=d\sqrt{T}.$
-[…]
-5 Improving the Estimation with Repeated Observations and Diversity
+Further, by a result in [Shamir, 2013], the tight regret bound for this problem is $\Theta(d\sqrt{T})$.
+Proposition 3.
+For any (possibly randomized) strategy the principal can adopt for posting contracts, there exists some parameter $\theta^{*}$ such that the term $\tt Reg$ can be lower bounded by $\Omega(d\sqrt{T})$.
+7 Improving Estimation and Learning with Repeated Observations and Diversity
 Algorithm 2 Use Repeated Observation as IV: Pure Exploitation with Diversity
 Parameters include time horizon $T$, number of tasks $d$, minimum eigenvalue $\lambda_{0}$, failure probability $\delta$
-Epoch $e$ contains |$\absolutevalue{\tau_{e}}=\max\{d\cdot 2^{e},8dC_{0}K_{0}\ln(d/\delta)/(c_{0}^{2}\lambda_{0}),4\sigma_{0}^{2}d^{2}\ln(d^{2}/\delta)/(c_{0}^{2}\lambda_{0})\}$ rounds
+Epoch $e$ contains |$\absolutevalue{\tau_{e}}=\max\{d\cdot 2^{e},8K_{0}\ln(d/\delta)/(\lambda_{0}),4\sigma_{0}^{2}d^{2}\ln(d^{2}/\delta)/(\lambda_{0})\}$ rounds
 In the 0-th epoch with $d$ rounds, post $\beta=\mathbf{e}_{i}$ for each $i\in[d]$ once and record $\widetilde{X}_{1},X_{1},Y_{1}$
@@
 end for
-This section continues the discussion where the principal must learn $\theta$ using sequential observations.
-The previous section showed that using an instrumental regression approach where the contract is the instrumental variable, the principal must balance exploration (ensuring that $\sigma_{\min}(B_{t}^{\top}X_{t})$ grows as $t$ increases) and exploitation (ensuring $\beta_{t}$ is close to the true optimal contract).
-In this section, I show that, under some additional assumptions, the learning rate of the principal can be greatly improved.
-Informally, I show that when the agents are sufficiently ‘diverse’, and that when repeated observations are available, the principal can achieve a logarithmic regret, even if the principal is using a ‘pure exploitation algorithm’.
-To state the results in a more concise manner, this section restricts attention to the case where the cost functions of the agents have diagonal quadratic forms.
-Specifically, the cost function of the agent is parameterized by a vector $\kappa\in(\mathbb{R}^{d})^{+}$, so that the cost function of the agent is $c(w)=\frac{1}{2}w^{\top}\mathtt{diag}(\kappa)^{-1}w$.
-Then, the agent’s best response to a contract $\beta$ is the maximizer of $\langle\beta,w\rangle-c(w)$, which is equal to
-$w=\kappa\odot\beta.$
-I assume that the agent cost function in each round, parameterized by $\kappa_{t}$, is drawn from an unknown distribution. The distribution satisfies the following.
+This section continues the discussion from the previous section, where the principal must learn or estimate $\theta^{*}$ using observational data.
+The previous section showed that using an instrumental regression approach and using the contract as the instrumental variable, the principal must balance exploration (ensuring that $\sigma_{\min}(B_{t}^{\top}X_{t})$ grows as $t$ increases) and exploitation (ensuring $\beta_{t}$ is close to the true optimal contract).
+In this section, we show that, under some additional assumptions, the learning rate of the principal can be greatly improved.
+Informally, we show that when the agents are sufficiently ‘diverse’, and that when repeated observations are available, the principal can achieve a logarithmic regret, even if the principal is using a ‘pure exploitation’ algorithm.
+[…]
+Maybe clash between $B_{t}$ and $B$.
+In addition to the boundedness of the the possible contract space $B$ in the previous section, we make two extra critical assumptions in this section.
+The first is that repeated signals are available.
+The second is that agents are drawn from some stochastic distribution and are sufficiently diverse.
+In the following, let us suppose that agents are drawn from some distribution $\mathcal{D}$, each element in the support of $D$ represents the cost function of the agent.
+It is assumed that for all elements in the support of $D$, the cost function is homogeneous of degree $k$ (i.e., it satisfies 1).
+Our first part of the assumption below states that agents are sufficiently diverse.
 Assumption 5.
-The minimum eigenvalue of $\mathbb{E}_{\kappa\sim\mathcal{D}}[\kappa\kappa^{\top}]$ is lower bounded by $\lambda_{0}$. Further, for any $\kappa$ in the support of $\mathcal{D}$, $\kappa\leq K_{0}$.
-The key part in the above assumption is the minimum eigenvalue $\lambda_{0}$, it suggests that the agents’ are sufficiently “diverse" in their types.
-The second assumption is that repeated observations are available.
-I assume that for each agent, the principal can observe two sets of signals, denoted by $x_{t}$ and $\widetilde{x}_{t}$.
+Repeated observations are available.
+We assume that in each round, the principal can observe two sets of signals, denoted by $x_{t}$ and $\widetilde{x}_{t}$.
 The signals are $\sigma_{0}^{2}$-subgaussian vectors and satisfy the following:
-$\mathbb{E}[x_{t}|w_{t}]=\mathbb{E}[\widetilde{x}_{t}|w_{t}]=w_{t}.$
+$\mathbb{E}[x_{t}|a_{t}]=\mathbb{E}[\widetilde{x}_{t}|a_{t}]=a_{t}.$
+They are conditionally independent given the agent’s hidden effort $a_{t}$:
+$x_{t}\perp\widetilde{x}_{t}|a_{t}.$
+The causal relationship between variables are summarized in Figure 1, with the dashed items included.
 Assumption 6.
-The two sets of signals are assumed to be (conditionally) independent. Specifically:
-$x_{t}\perp\widetilde{x}_{t}|w_{t}.$
+Agents are sufficiently diverse in their talent. Fix any $\beta\in\mathcal{B}$, we have that the following holds
+$\lambda_{\min}\mathbb{E}_{c(\cdot)\sim\mathcal{D}}[a(\beta;c(\cdot))^{\top}a(\beta;c(\cdot))]\geq\lambda_{0}.$
+The first part of the assumption requires repeated observations, which is often realistic in practical scenarios.
+Some literature in economics also study the problem of repeated outputs (though under very different contexts, for more discussion see Chapter 10.2 in [Bolton and Dewatripont, 2004]).
+The second part of the assumptions is essentially requireing that there is sufficient diverity in the agent’s ability to complete tasks. We give a concrete example satisfying assumption below.
+Example 1.
+Consider the cost functions of the agents have diagonal quadratic forms.
+Specifically, the cost function of the agent is parameterized by a vector $\kappa\in(\mathbb{R}^{d})^{+}$, so that the cost function of the agent is $c(a)=\frac{1}{2}a^{\top}\mathtt{diag}(\kappa)^{-1}a$.
+Then, the agent’s best response to a contract $\beta$ is the maximizer of $\langle\beta,a\rangle-c(a)$, which is equal to
+$a=\kappa\odot\beta.$
+We assume that the agent cost function in each round, parameterized by $\kappa_{t}$, is drawn from an unknown distribution. The distribution satisfies the following.
+$\lambda_{\min}\mathbb{E}_{\kappa\sim\mathcal{D}}[\kappa\kappa^{\top}]\geq\lambda.$
+Then we know that
+$\lambda_{\min}\mathbb{E}[a(\beta)a^{\top}(\beta)]=c_{0}^{2}\cdot\lambda.$
+Hence the diversity condition is satisfied. Note that this model has also appeared in the prior work [Thiele, 2010].
+In addition to the above two assumptions, we require a relatively mild boundedness assumption as below.
+Assumption 7.
+Fix any $\beta\in B$. We have the following for any agent type: $\norm{a}_{2}^{2}\leq K_{0}$. Here $a$ is the best response to contract $\beta$.
+7.1 Offline Setting: An Alternative GMM Estimator
 The causal relationship is summarized in Figure 1 with the dashed items included.
-It can be observed that the second set of observations can serve as an instrumental variable for the other, specifically:
+It can be observed that the second set of observations can serve as an instrumental variable for the other, and that we have:
 $\mathbb{E}[\widetilde{x}_{t}(y_{t}-\langle x_{t},\theta\rangle)]=\mathbf{0}\ (\in\mathbb{R}^{d}).$
-Therefore, we can use the following GMM estimator to obtain an estimate $\hat{\theta}$:
+Assume the principal is given some offline data as $((\beta_{i},x_{i},\widetilde{x}_{i},y_{i})_{i=1}^{T})$.
+Then, the principal can use the following GMM estimator to obtain an estimate $\hat{\theta}$:
 $\hat{\theta}_{T}=(\widetilde{X}_{T}^{\top}X_{T})^{-1}\widetilde{X}_{T}^{\top}Y_{T}.$
+Here $X_{T},\widetilde{X}_{T}$ are a $T\times d$ matrices where row $t$ is $x_{t}$ or $\widetilde{x}_{t}$ respectively; $Y_{T}$ is a vector with the $t$-th entry being $y_{t}$.
+We state the convergence result for the GMM estimator above.
+Proposition 4.
+With probability $1-\delta$,
+$\norm{\theta-\theta^{*}}_{2}\leq\frac{\sqrt{dT\log(dT/\delta)}}{\sigma_{\min}(\widetilde{X}_{T}^{\top}X_{T})}.$
+The proof is straightforward and similar to that of Proposition 1.
+7.2 Online Setting: Pure exploitation algorithm
+We again turn to the online setting where the principal must learn the optimal contract.
 The algorithm will simply use the estimator above, without regard to ‘exploration’.
@@
 Remark 4.
-At a technical level, the analysis is a combination of the convergence analysis of GMM estimators (e.g. Chapter 18 in [5]) and smoothed analysis in bandit problems ([19, 20, 21]).
+At a technical level, the analysis is a combination of the convergence analysis of GMM estimators (e.g.
+Chapter 18 in [Greene, 2008]) and smoothed analysis in bandit problems ([Bastani et al., 2021, Kannan et al., 2018, Sivakumar et al., 2022]).
+8 Proof of Robustness Result in Theorem 1
+We will show that for any contract $w$, there exists an non-negative affine contract $\phi$ and compatible distributions $\mathcal{F}$ such that
+$u(\phi;\mathcal{F})\geq u(w;\mathcal{F}).$
+Since under an affine contract, the agent’s expected payment for any action $a$ remains the same for any distribution, we know
+$u(\phi)=u(\phi;\mathcal{F}).$
+Moreover, by definition
+$u(w)\leq u(w;\mathcal{F}).$
+By the above we know that
+$u(\beta)\geq u(w).$
+Then using Lemma 1, which shows we can further improve an affine contract to a linear contract, the theorem is proven.
+We define the following set of points:
+[…]
+For each upper hyperplane $H$ supporting the set $S$ from above, consider the set of points that the plane $H$ supports, and project these points back to the space $\mathcal{X}$.
+Denote this set as $P(H)$.
+By definition $P(H)\subset\mathcal{X}$.
+By a classic result in convex geometry (e.g.
+[Lee and Santos, 2017], [De Loera et al., 2010]), if we consider the set of upper facets of the set $S$ (for which each defines an upper hyperplane), projecting each facet back to $\mathcal{X}$ in fact gives an exact partition of $\mathcal{X}$, i.e.,
+$\cup_{H:H\text{ touches an upper facet}}\mathtt{conv}(P(H))=\mathtt{conv}(\mathcal{X}).$
+For our purposes, we will be considering upper supporting hyperplanes whose normal vector’s last component is bounded away from 0.
+lower bounded by some $\varepsilon$.
+In the following denote $n_{d+1}(H)$ to as the last component of the normal vector.
+For hyperplanes with $n_{d+1}(H)=0$, they do not represent valid contracts.
+Lemma 2.
+[…]
+$\cup_{H\in\mathcal{H}}\mathtt{conv}(P(H))=\mathtt{conv}(\mathcal{X}).$
+Proof.
+The set $S$ has at most a finite number of upper facets, to which each corresponds to a hyperplane $H$.
+Then, simply set $\varepsilon$ to be the minimum of $n_{d+1}(H)$ among all such $H$, and the lemma is proven.
+∎
+Now, it can be seen that the set $\mathcal{H}$ as in the previous lemma is topologically equivalent to a closed and convex set.
+This will allow us to apply fixed-point theorems in the subsequent analysis.
+In the following, we define $\phi(H)$ to be the affine contract associated with the hyperplane $H$, and use $a(H)=a(\phi(H))$ to denote the action that is induced by $\phi(H)$.
+Lemma 3.
+There exists an upper supporting hyperplane $H$, such that the incentivized action $a^{*}$ under $\phi(H)$ lies in the set $\mathtt{conv}(P(H))$.
+Recall this is the notion of “self-owned” hyperplanes as introduced previously.
+Proof.
+By the previous lemma, the union of all $\mathtt{conv}(P(H))$ cover the entire space $\mathcal{A}$.
+Then for any $a$, there must exist some hyperplane $H^{\prime}$, such that $a(H)\in\mathtt{conv}(P(H^{\prime}))$.
+So, let us define a function $g:\mathcal{H}\rightarrow 2^{\mathcal{H}}$, such that for each upper supporting hyperplane $H$, the function maps to a set of supporting hyperplanes defined as follows:
+[…]
+By the previous lemma, the set $g(H)$ is non-empty.
+Further, we can see that the set $g(H)$ is convex and that $g$ has a closed graph (see Section A.3 for a brief discussion).
+By Katukani’s fixed point theorem, there exists some hyperplane $H$ such that $g(H)=H$.
+Hence the theorem is proven.
+∎
+We can now show that there exists some affine contract which does no worse than $w$ in worse-case.
+Lemma 4.
+Let $H$ be the hyperplane as in the statement of the previous lemma (i.e., “self-owned”) and $H$ induces $a(H)$.
+Then there exists some distribution under which the contract $w$ also induces $a(H)$.
+Further, $\phi(H)$ uses no greater payment than $w(\cdot)$.
+Proof.
+First note that under any distribution, for any action $a\in\mathcal{A}$, the payment to the agent under $w$ is certainly no larger than under $\phi$, since $\phi$ bounds $w$ from above.
+Now, since $H$ is “self-owned”, we have $a(H)\in\mathtt{conv}(P(H))$.
+We can define a compatible distribution where the signal distribution of action $a(H)$ is supported only on points in $P(H)$.
+For points $x\in P(H)$ we have $\phi(x)=w(x)$.
+Then, the payment to the agent if he takes action $a(H)$ is exactly the same under $\phi(H)$ as under $w(\cdot)$.
+Hence, under this distribution, the contract $w$ also induces action $a(\beta)$.
+∎
+Thus so far, we have shown that any contract $w(\cdot)$ can be improved to an affine contract whose worst-case payoff is no less than that of $w(\cdot)$.
+By Lemma 1, any affine contract can be improved to a linear contract.
+This completes the proof.
+Our main result in this section Theorem 1 then follows from the above lemmas.
 Remark 5.
-Interestingly, the “fast" convergence also holds under a different notion of ‘diversity’.
-Instead of assuming agents’ cost functions are diverse, consider a setting where agents do not necessarily best respond, and their response $w_{t}$ may be a random vector with some variance lower bounded from below.
-In this setting, the minimum singular value $\widetilde{X}_{t}^{\top}X_{t}$ still grows linearly with $t$, and the “fast" convergence rate is achieved.
-6 Numerical Experiments
-Figure 2: (a): Total loss using the Contract as IV, and exploration is achieved through small random perturbations; (b): Total loss using the repeated observations as IV and when the “diversity" condition is satisfied.
-This section describes numerical simulations that test the performance of the learning algorithms based on the GMM estimators.
-In the experiments, the number of tasks is chosen as $d=5$, the total number of rounds is chosen as $T=10^{6}$, and the true parameter $\theta$ is chosen so that each entry is between $[1,3]$.
-Experiments are ran on a personal laptop.
-The experiments have two settings.
-In the first setting, the contract is used as the instrumental variable, and small random perturbations are added as a means to achieve exploration in the exploration phase (Algorithm 1).
-The small perturbations are chosen as truncated Gaussian vectors so that the posted contract has components within $[1,3]$.
-The agents have cost function $c(w)=\norm{w}_{2}^{2}$.
-The cumulative utility loss of a single run is plotted in Figure 2(a).
-It can be seen the loss grows quickly in the warm start phase (which consists of $d\sqrt{T}\approx 5\times 10^{3}$ rounds), and then the regret grows slowly in the remaining rounds.
-The experiments are run for 20 trials, and the average cumulative loss in the 20 trials is 4134.7, roughly on the order of $O(d\sqrt{T})$.
-In the second setting, repeated observations are used as the instrumental variable (Algorithm 2). The agent’s cost function is chosen as a diagonal quadratic function with form
-$c(w)=\sum_{i=1}^{d}\frac{w_{i}^{2}}{2k_{i}}.$
-The parameter $k_{i}$ is sampled from the uniform distribution on $[1,2]$ so that the ‘diversity’ condition is satisfied.
-The cumulative loss of a single run is plotted in Figure 2(b).
-The loss grows quickly in the 1st epoch (consisting of roughly $\widetilde{O}(d)$ rounds) and then grows very slowly in the remaining rounds.
-The experiment is repeated for 20 trials, and the average cumulative loss is 89.6 (recall the theoretical limit is $\widetilde{O}(d)$).
-7 Conclusion and Limitation
-This work studied the multitask principal-agent problem.
-I first present a result justifying the use of uniform contracts when the principal contracts with heterogeneous agents.
-Then, I study the problem of learning the optimal contract and identify the problem as a measurement error problem, and propose using instrumental regression with the generalized method of moments estimator to estimate and learn the optimal contract.
-A limitation of the current work is that in the “diversity” condition in Section 5, the agents are limited to have quadratic cost functions.
-I believe this can be extended to more general cost functions (beyond quadratic functions and beyond $k=2$).
-Specifically, the “diversity” condition will hold whenever there are some variations in the agent’s “abilities" in each task.
-I will leave a more formal treatment to future work.
-Appendix A Proof of Theorem 2
+Our results can also be extended to the case when the signal space is infinite, e.g., $\mathcal{X}=[0,1]^{d}$.
+However, this requires some additional regularity assumptions.
+In particular, we need to ensure the contract does not grow “too quickly” at the boundary.
+This avoids the case that at the boundary a “vertical” supporting hyperplane occurs, i.e., one with $n_{d+1}(H)=0$.
+Such hyperplanes does not correspond to valid contracts, and Lemma 2 does not hold.
+9 Proof of Regret in Theorem 3
 Recall the estimator in each epoch
 $\hat{\theta}_{e}=\left(\widetilde{X}_{e}^{\top}X_{e}\right)^{-1}\widetilde{X}_{e}Y_{e}.$
-Denote $\gamma_{t}=y_{t}-\langle x_{t},\theta\rangle$, and collect $\gamma_{t}$ into the column vector $\Gamma_{e}$.
-Proposition 2.
-$\norm{\hat{\theta}_{e}-\theta}_{2}\leq\frac{\norm{\widetilde{X}_{e}\Gamma_{e}}}{\sigma_{\min}(\widetilde{X}_{e}^{\top}X_{e})}$.
+Denote $\gamma_{t}=y_{t}-\langle x_{t},\theta\rangle$, and collect $\gamma_{t}$ into the column vector $\Gamma_{e}$. Recall that by Proposition 4, we have
+$\norm{\hat{\theta}_{e}-\theta}_{2}\leq\frac{\norm{\widetilde{X}_{e}\Gamma_{e}}}{\sigma_{\min}(\widetilde{X}_{e}^{\top}X_{e})}.$ (2)
 To bound the estimation error, the following will lower bound $\sigma_{\min}(\widetilde{X}_{e}^{\top}X_{e})$ and upper bound $\norm{\widetilde{X}_{e}\Gamma_{e}}$ separately.
@@
 Lemma 6.
-With high probability $\delta$, $\lambda_{\min}(\sum_{t\in\tau_{e}}w_{t}w_{t}^{\top})\geq\absolutevalue{\tau_{e}}c_{0}^{2}\lambda_{0}/2$.
+With high probability $\delta$, $\lambda_{\min}(\sum_{t\in\tau_{e}}a_{t}a_{t}^{\top})\geq\absolutevalue{\tau_{e}}\lambda_{0}/2$.
 Proof.
 […]
-$\mathbb{E}[W_{t}]$ $=\mathbb{E}[w_{t}w_{t}^{\top}]$
-$=\mathbb{E}[\mathtt{diag}(\kappa_{t})\theta_{t}\theta_{t}^{\top}\mathtt{diag}(\kappa_{t})]$
-$=\mathbb{E}[\Theta_{t}\kappa_{t}\kappa_{t}^{\top}\Theta_{t}]$
-$=\Theta_{t}\mathbb{E}[\kappa_{t}\kappa_{t}^{\top}]\Theta_{t}$
-Recall $\lambda_{\min}(\mathbb{E}[\kappa_{t}\kappa_{t}^{\top}])\geq\lambda_{0}$. The following holds
-$\lambda_{\min}(\mathbb{E}W_{t})\geq c_{0}^{2}\lambda_{0}.$
+By the diversity condition,
+$\lambda_{\min}(\mathbb{E}W_{t})\geq\lambda_{0}.$
 By Tropp’s lemma,
-$\Pr[\lambda_{\min}(\sum_{k}W_{k})\leq\absolutevalue{\tau_{e}}c_{0}^{2}\lambda_{0}/2]$ $\leq d\cdot\exp(-\absolutevalue{\tau_{e}}c_{0}^{2}\lambda_{0}/(8dC_{0}K_{0}))$
+$\Pr[\lambda_{\min}(\sum_{k}W_{k})\leq\absolutevalue{\tau_{e}}\lambda_{0}/2]$ $\leq d\cdot\exp(-\absolutevalue{\tau_{e}}\lambda_{0}/(8K_{0}))$
 $\leq\delta$
 when
-|$\absolutevalue{\tau_{e}}\geq\frac{8dC_{0}K_{0}}{c_{0}^{2}\lambda_{0}}\cdot\ln(d/\delta)$
+|$\absolutevalue{\tau_{e}}\geq\frac{8K_{0}}{\lambda_{0}}\cdot\ln(d/\delta)$
 Lemma 7.
-With probability $1-2\delta$, for each epoch $e$ with length as specified in Algorithm 2, $\sigma_{\min}(\widetilde{X}_{e}^{\top}X_{e})\geq\absolutevalue{\tau_{e}}c_{0}^{2}\lambda_{0}/4$.
+With probability $1-2\delta$, for each epoch $e$ with length as specified in Algorithm 2, $\sigma_{\min}(\widetilde{X}_{e}^{\top}X_{e})\geq\absolutevalue{\tau_{e}}\lambda_{0}/4$.
 Proof.
-In the following denote $\varepsilon_{t}=x_{t}-w_{t},\widetilde{\varepsilon}_{t}=\widetilde{x}_{t}-w_{t}$.
+In the following denote $\varepsilon_{t}=x_{t}-a_{t},\widetilde{\varepsilon}_{t}=\widetilde{x}_{t}-a_{t}$.
 $\widetilde{X}_{e}^{\top}X_{e}$ $=\sum_{t\in\tau_{e}}x_{t}\widetilde{x}_{t}^{\top}$
-$=\sum_{t\in\tau_{e}}(w_{t}+\varepsilon_{t})(w_{t}+\widetilde{\varepsilon}_{t})^{\top}$
-$=\sum_{t\in\tau_{e}}w_{t}^{\top}w_{t}+\varepsilon_{t}^{\top}w_{t}+w_{t}\varepsilon_{t}+\varepsilon_{t}^{\top}\widetilde{\varepsilon}_{t}$
+$=\sum_{t\in\tau_{e}}(a_{t}+\varepsilon_{t})(a_{t}+\widetilde{\varepsilon}_{t})^{\top}$
+$=\sum_{t\in\tau_{e}}a_{t}^{\top}a_{t}+\varepsilon_{t}^{\top}a_{t}+a_{t}\varepsilon_{t}+\varepsilon_{t}^{\top}\widetilde{\varepsilon}_{t}$
 The previous lemma showed that with probability $1-\delta$,
-$\lambda_{\min}(\sum_{t\in\tau_{e}}w_{t}^{\top}w_{t})\geq\absolutevalue{\tau_{e}}c_{0}^{2}\lambda_{0}/2.$
+$\lambda_{\min}(\sum_{t\in\tau_{e}}a_{t}^{\top}a_{t})\geq\absolutevalue{\tau_{e}}\lambda_{0}/2.$
 The remaining term
-$\sum_{t\in\tau_{e}}\varepsilon_{t}w_{t}^{\top}+w_{t}\widetilde{\varepsilon}_{t}^{\top}+\varepsilon\widetilde{\varepsilon}^{\top}$
+$\sum_{t\in\tau_{e}}\varepsilon_{t}a_{t}^{\top}+a_{t}\widetilde{\varepsilon}_{t}^{\top}+\varepsilon\widetilde{\varepsilon}^{\top}$
 is a random matrix with mean $\mathbf{0}$.
@@
 […]
-|$\geq\absolutevalue{\tau_{e}}c_{0}^{2}\lambda_{0}/4.$
+|$\geq\absolutevalue{\tau_{e}}\lambda_{0}/4.$
 […]
@@
 Recall $\gamma_{t}=y_{t}-\langle x_{t},\theta\rangle$, and $\Gamma_{e}$ is the column vector collecting the variables $\gamma_{t}$.
-Note that $\gamma_{t}\widetilde{x}_{t}$ is a zero-mean, $\sigma_{0}C_{0}K_{0}$-subgaussian random vector and independent conditioned on all previous steps.
-The result then follows from a standard concentration inequality.
+Note that $\gamma_{t}\widetilde{x}_{t}$ is a zero-mean, $\sigma_{0}K_{0}$-subgaussian random vector and independent conditioned on all previous steps.
+The result then follows from the standard concentration inequality (see Section A.1).
 ∎
@@
 Proof.
-By Lemma 4, Lemma 3, and Proposition 2, with probability $1-3\delta$, the regret in each epoch can be upper bounded as |$\absolutevalue{\tau_{e}}\cdot\widetilde{O}(d/(\absolutevalue{\tau_{e}}\lambda_{0}^{2}))=\widetilde{O}(d/\lambda_{0}^{2})$.
+By Lemma 8, Lemma 7, and Equation 2, with probability $1-3\delta$, the regret in each epoch can be upper bounded as |$\absolutevalue{\tau_{e}}\cdot\widetilde{O}(d/(\absolutevalue{\tau_{e}}\lambda_{0}^{2}))=\widetilde{O}(d/\lambda_{0}^{2})$.
 Choose some appropriate $\delta$, e.g., $\delta<O(1/T^{2})$.
@@
 ∎
+10 Conclusion
+In this work, we examined the role of linear contracts in principal-agent problems from three key perspectives: robustness, uniformity, and learning.
+Firstly, we analyzed the robustness properties of linear contracts.
+In an ambiguous principal-agent setting where only first-moment information is available, we demonstrated that a linear contract is worst-case optimal.
+This result underscores the resilience of linear contracts under uncertainty, providing a rationale for their widespread use in practical applications.
+A natural question for future research is to investigate the extent to which our assumptions can be relaxed while preserving the robustness guarantees.
+Understanding the boundary conditions of these results could yield deeper insights into the trade-offs between contract simplicity and performance in uncertain environments.
+Second, we provided a theoretical justification for the use of uniform contracts when contracting with heterogeneous agents.
+Our analysis establishes conditions under which uniform contracts are not only feasible but also optimal.
+Specifically, when agents exhibit certain cost function properties, a uniform contract structure can align incentives effectively without the need for agent-specific tailoring.
+This result is particularly relevant in settings such as franchising or platform-based economies, where principals interact with diverse agents yet often employ standardized contracts.
+Finally, we investigated the problem of learning optimal contracts in scenarios where the principal lacks complete knowledge of agent-specific parameters.
+We drew a connection between contract learning and instrumental regression, highlighting how techniques from causal inference can inform strategic contract design.
+Notably, recent research has recognized the relationship between learning in strategic settings and causal inference methods, such as instrumental regression.
+Building on this, a promising direction for future work is to explore how learning with strategic interactions in Stackelberg games can be framed within the instrumental regression framework.
+This could open new avenues for leveraging econometric techniques in the study of contract optimization.
+Appendix A Appendix
+A.1 Concentration Lemma
+Lemma 9.
+Let $\gamma_{1},\dots,\gamma_{T}$ be independent $\sigma$-subgaussian random variables.
+Let $v_{1},\dots,v_{T}$ be vectors in $\mathbb{R}^{d}$ with each $v_{T}$ chosen arbitrarily as a function of $(v_{1},\gamma_{1}),\dots,(v_{t-1},\gamma_{t-1})$ subject to $\norm{v_{t}}\leq C$.
+Then with probability at least $1-\delta$,
+$\norm{\sum_{t=1}^{T}\gamma_{t}v_{t}}\leq\sqrt{2dC\sigma T\log(Td/\delta)}.$
+A.2 Proof for Proposition 2
+Assumption 8.
+There exists some constant $\underline{a}>0$, such that under the contract $\beta=e_{i}$, any agent’s private effort satisfies $a_{i}>\underline{a}$.
+Proof of Proposition 2.
+Fix a dimension (task) $i\in[d]$.
+Consider the $O(\sqrt{T})$ rounds in which the principal posted the contract $\beta=\mathbf{e}_{i}$.
+Let the sum of the signal for the $i$-th task be $X_{i}$ over these rounds, and let the sum of agents’ true effort be $A_{i}$.
+Note that in the denominator of the GMM estimator, the matrix $B_{\tau}X_{\tau}$ simply becomes $\mathtt{diag}(X_{1},\dots,X_{d})$. Using proposition 1, we have with probability $1-\delta$,
+$\norm{\hat{\theta}-\theta}_{2}\leq\frac{\sqrt{d\sqrt{T}\log(1/\delta)}}{\min(X_{i})}.$
+Hence let us now turn to bound $\min(X_{i})$. Note that with probability $1-\delta_{0}$, by a standard Hoeffding inequality, we have
+$\Pr[\absolutevalue{X_{i}-W_{i}}<O\left(\sqrt{\frac{\sigma^{2}\ln(1/\delta_{0})}{T^{1/2}}}\right)]<\delta_{0}.$
+Hence choosing $\delta_{0}=\delta/d$, with probability at least $1-\delta$, for large enough $T$, we have $\min_{i}X_{i}\geq\Omega(T^{1/2})$.
+In the initial exploration rounds, the estimation error after the initial $d\sqrt{T}$ rounds is:
+$\norm{\hat{\theta}-\theta^{*}}_{2}=O(\sqrt{d\log(1/\delta)/\sqrt{T}}).$
+The regret in subsequent rounds can be bounded as:
+$T\cdot\norm{\hat{\theta}-\theta^{*}}_{2}^{2}=O(d\sqrt{T}\cdot\log(1/\delta)).$
+Hence the total regret is $O(d\sqrt{T}\cdot\log(1/\delta))$. ∎
+A.3 Application of Kakutani’s Fixed-Point Theorem
+Applying Kakutani’s fixed-point theorem requires the following conditions on the mapping $g:\mathcal{H}\rightarrow 2^{\mathcal{H}}$.
+For any $x$, $g(H)$ is non-empty, convex, and closed.
+Further $\phi$ has is upper hemicontinuous (which together with closeness of $g(H)$ implies $g$ have a close graph).
+One of the key conditions, non-emptiness, is guaranteed by Lemma 2.
+To see that $g(H)$ is convex, note that if two hyperplanes $H_{1}$ and $H_{2}$ supports $a(H)$, so does their convex combination.
+To see $g(H)$ is closed, note that for any sequence of hyperplanes $(H^{\prime}_{i})$ which converge to $H^{\prime}$, if all hyperplanes in the sequence $H^{\prime}_{i}$ supports $a(H)$, then so does the limit $H^{\prime}$.
+To show upper hemicontinuity, let $(H^{\prime}_{i})$ be a sequence of hyperplanes that converge to $H^{\prime}$, and let $(H_{i})$ be a sequence of points that converge to $H$.
+Note that if for any $i$ we have that $H^{\prime}_{i}$ supports $a(H_{i})$, then in the limit $H^{\prime}$ supports $a(H)$.
+Hence upper hemicontinuity is shown.